### 01 Feb 2024 v1

Abstract
[…]
Supersymmetry, Higgs boson decay, CP violation
pacs:
12.60.Jv, 14.80.Da
I Introduction
In 2012, the 125 GeV Higgs boson discovered by the Large Hadron Collider (LHC) mh-ATLAS ; mh-CMS , the measured mass of the Higgs boson now is PDG1 $m_{h}=125.25\pm 0.17\>{\rm{GeV}}.$ The discovery of the Higgs boson marked a huge success for the Standard Model (SM), but it did not stop the search for New Physics (NP) at the LHC, and one of them was the search for lighter scalar particles.
[…]
As one of the extensions of the SM, the $\mu$ from $\nu$ supersymmetric standard model ($\mu$$\nu$SSM) mnSSM ; mnSSM1 ; mnSSM1-1 ; parameter-space ; mnSSM2 ; mnSSM2-1 ; Zhang1 ; Zhang2 can solve the $\mu$ problem m-problem of the minimal supersymmetric standard model (MSSM) MSSM ; MSSM1 ; MSSM2 ; MSSM3 ; MSSM4 , through introducing three singlet right-handed neutrino superfields $\hat{\nu}_{i}^{c}$ ($i=1,2,3$). The neutrino superfields lead the mixing of the neutral components of the Higgs doublets with the right-handed sneutrinos, that is different from the Higgs sector of the MSSM. The mixing can change the Higgs couplings, and influence the decay processes of the Higgs bosons. In addition, we also introduce CP-violation, and we also get a lighter Higgs at $\sim$ 95 GeV with a suitable parameter space.
The paper is organized as follows. In Sec. II, we introduce the CPV $\mu\nu$SSM briefly, about the superpotential and the CPV phases. In Sec. III, we study the excess at 95 GeV in the CPV $\mu\nu$SSM. Sec. IV and Sec. V, we show the numerical analysis and the conclusion.
II The Model
The superpotential of the $\mu\nu$SSM contains Yukawa couplings for neutrinos, two additional types of terms involving the Higgs doublet superfields $\hat{H}_{u}$ and $\hat{H}_{d}$, and the right-handed neutrino superfields $\hat{\nu}_{i}^{c}$mnSSM
$W={\epsilon_{ab}}\left({Y_{{u_{ij}}}}\hat{H}_{u}^{b}\hat{Q}_{i}^{a}\hat{u}_{j}^{c}+{Y_{{d_{ij}}}}\hat{H}_{d}^{a}\hat{Q}_{i}^{b}\hat{d}_{j}^{c}+{Y_{{e_{ij}}}}\hat{H}_{d}^{a}\hat{L}_{i}^{b}\hat{e}_{j}^{c}\right)$
$\hskip 27.03003pt+{\epsilon_{ab}}{Y_{{\nu_{ij}}}}\hat{H}_{u}^{b}\hat{L}_{i}^{a}\hat{\nu}_{j}^{c}-{\epsilon_{ab}}{\lambda_{i}}e^{i\phi_{\lambda_{i}}}\hat{\nu}_{i}^{c}\hat{H}_{d}^{a}\hat{H}_{u}^{b}+\frac{1}{3}{\kappa_{ijk}}\hat{\nu}_{i}^{c}\hat{\nu}_{j}^{c}\hat{\nu}_{k}^{c},$ (1)
where $\hat{H}_{u}^{T}=\Big{(}{\hat{H}_{u}^{+},\hat{H}_{u}^{0}}\Big{)}$, $\hat{H}_{d}^{T}=\Big{(}{\hat{H}_{d}^{0},\hat{H}_{d}^{-}}\Big{)}$, $\hat{Q}_{i}^{T}=\Big{(}{{{\hat{u}}_{i}},{{\hat{d}}_{i}}}\Big{)}$, $\hat{L}_{i}^{T}=\Big{(}{{{\hat{\nu}}_{i}},{{\hat{e}}_{i}}}\Big{)}$ (the index $T$ denotes the transposition) represent the MSSM-like doublet Higgs superfields, and $\hat{u}_{i}^{c}$, $\hat{d}_{i}^{c}$, and $\hat{e}_{i}^{c}$ are the singlet up-type quark, down-type quark and charged lepton superfields, respectively. In addition, $Y_{u,d,e,\nu}$, $\lambda$, and $\kappa$ are dimensionless matrices, a vector, and a totally symmetric tensor. $a,b=1,2$ are SU(2) indices with antisymmetric tensor $\epsilon_{12}=1$, and $i,j,k=1,2,3$ are generation indices. The CP is violated by the parameter $\lambda_{i}$, the CP-violating phase is $\phi_{\lambda_{i}}$.
In the superpotential, if the scalar potential is such that nonzero vacuum expectative values (VEVs) of the scalar components ($\tilde{\nu}_{i}^{c}$) of the singlet neutrino superfields $\hat{\nu}_{i}^{c}$ are induced, the effective bilinear terms $\epsilon_{ab}\varepsilon_{i}\hat{H}_{u}^{b}\hat{L}_{i}^{a}$ and $\epsilon_{ab}\mu\hat{H}_{d}^{a}\hat{H}_{u}^{b}$ are generated, with $\varepsilon_{i}=Y_{\nu_{ij}}\left\langle{\tilde{\nu}_{j}^{c}}\right\rangle$ and $\mu={\lambda_{i}}\left\langle{\tilde{\nu}_{i}^{c}}\right\rangle$, once the electroweak symmetry is broken. The last term Eq. (1) generates the effective Majorana masses for neutrinos at the electroweak scale. Therefore, the $\mu\nu$SSM can generate three tiny neutrino masses at the tree level through TeV scale seesaw mechanism mnSSM ; neutrino-mass ; neu-mass1 ; neu-mass2 ; neu-mass3 ; neu-mass4 ; neu-mass5 ; neu-mass6 .
It is worth explaining why TeV scale seesaw was chosen. Through a seesaw on the scale of the Grand Unified Theory (GUT), one can get the Yukawa couplings of order one for neutrinos. But we know that the Yukawa coupling of the electron is on the order of $10^{-6}$, and the Yukawa couplings of neutrinos can also be around on the order of $10^{-6}$ instead of one. In the TeV scale seesaw, this is sufficient to reproduce the neutrino mass, if the Yukawa coupling of the neutrino is of the same order as the Yukawa coupling of the electron mnSSM . Here it is important to note that the VEVs of the left-handed sneutrinos $\upsilon_{\nu_{i}}$ are generally small. We know that the Dirac masses for the neutrinos $m_{D_{i}}=Y_{\upsilon_{\nu_{i}}}\upsilon_{u}\lesssim 10^{-4}$ GeV in the TeV scale seesaw. So we can get an estimate of the VEVs of the left-handed sneutrinos, $\upsilon_{\nu_{i}}\lesssim m_{D_{i}}\lesssim 10^{-4}$ GeV, which means that $\upsilon_{\nu_{i}}\ll\upsilon_{d},\upsilon_{u}$ mnSSM ; mnSSM1 .
The general soft SUSY-breaking terms of the $\mu\nu$SSM are given by
$-\mathcal{L}_{soft}=m_{{{\tilde{Q}}_{ij}}}^{\rm{2}}\tilde{Q}{{}_{i}^{a\ast}}\tilde{Q}_{j}^{a}+m_{\tilde{u}_{ij}^{c}}^{\rm{2}}\tilde{u}{{}_{i}^{c\ast}}\tilde{u}_{j}^{c}+m_{\tilde{d}_{ij}^{c}}^{2}\tilde{d}{{}_{i}^{c\ast}}\tilde{d}_{j}^{c}+m_{{{\tilde{L}}_{ij}}}^{2}\tilde{L}_{i}^{a\ast}\tilde{L}_{j}^{a}$
$\hskip 48.36958pt+m_{\tilde{e}_{ij}^{c}}^{2}\tilde{e}{{}_{i}^{c\ast}}\tilde{e}_{j}^{c}+m_{{H_{d}}}^{\rm{2}}H_{d}^{a\ast}H_{d}^{a}+m_{{H_{u}}}^{2}H{{}_{u}^{a\ast}}H_{u}^{a}+m_{\tilde{\nu}_{ij}^{c}}^{2}\tilde{\nu}{{}_{i}^{c\ast}}\tilde{\nu}_{j}^{c}$
$\hskip 48.36958pt+\epsilon_{ab}{\left[{{({A_{u}}{Y_{u}})}_{ij}}H_{u}^{b}\tilde{Q}_{i}^{a}\tilde{u}_{j}^{c}+{{({A_{d}}{Y_{d}})}_{ij}}H_{d}^{a}\tilde{Q}_{i}^{b}\tilde{d}_{j}^{c}+{{({A_{e}}{Y_{e}})}_{ij}}H_{d}^{a}\tilde{L}_{i}^{b}\tilde{e}_{j}^{c}+{\rm{H.c.}}\right]}$
$\hskip 48.36958pt+\left[{\epsilon_{ab}}{{({A_{\nu}}{Y_{\nu}})}_{ij}}H_{u}^{b}\tilde{L}_{i}^{a}\tilde{\nu}_{j}^{c}-{\epsilon_{ab}}{{({A_{\lambda}}e^{i\phi_{\lambda}}\lambda)}_{i}}\tilde{\nu}_{i}^{c}H_{d}^{a}H_{u}^{b}+\frac{1}{3}{{({A_{\kappa}}\kappa)}_{ijk}}\tilde{\nu}_{i}^{c}\tilde{\nu}_{j}^{c}\tilde{\nu}_{k}^{c}+{\rm{H.c.}}\right]$
$\hskip 48.36958pt-\frac{1}{2}\left({M_{3}}{{\tilde{\lambda}}_{3}}{{\tilde{\lambda}}_{3}}+{M_{2}}{{\tilde{\lambda}}_{2}}{{\tilde{\lambda}}_{2}}+{M_{1}}{{\tilde{\lambda}}_{1}}{{\tilde{\lambda}}_{1}}+{\rm{H.c.}}\right).$ (2)
Here, the first two lines contain mass squared terms of squarks, sleptons, and Higgses. The next two lines consist of the trilinear scalar couplings. In the last line, $M_{3}$, $M_{2}$, and $M_{1}$ denote Majorana masses corresponding to $SU(3)$, $SU(2)$, and $U(1)$ gauginos $\tilde{\lambda}_{3}$, $\tilde{\lambda}_{2}$, and $\tilde{\lambda}_{1}$, respectively. In addition to the terms from $\mathcal{L}_{soft}$, the tree-level scalar potential receives the usual $D$- and $F$-term contributions mnSSM1 ; mnSSM1-1 .
Once the electroweak symmetry is spontaneously broken, the neutral scalars develop in general the VEVs. The CP can violated by the VEVs of the scalar fields:
$\langle H_{d}^{0}\rangle=e^{i\phi_{\upsilon_{d}}}\upsilon_{d},\qquad\langle H_{u}^{0}\rangle=e^{i\phi_{\upsilon_{u}}}\upsilon_{u},\qquad\langle\tilde{\nu}_{i}\rangle=\upsilon_{\nu_{i}},\qquad\langle\tilde{\nu}_{i}^{c}\rangle=e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}\upsilon_{\nu_{i}^{c}}.$ (3)
One can define the neutral scalars as
$H_{d}^{0}=e^{i\phi_{\upsilon_{d}}}(\frac{h_{d}+iP_{d}}{\sqrt{2}}+\upsilon_{d}),\qquad\;\tilde{\nu}_{i}=\frac{(\tilde{\nu}_{i})^{\Re}+i(\tilde{\nu}_{i})^{\Im}}{\sqrt{2}}+\upsilon_{\nu_{i}},$
$H_{u}^{0}=e^{i\phi_{\upsilon_{u}}}(\frac{h_{u}+iP_{u}}{\sqrt{2}}+\upsilon_{u}),\qquad\tilde{\nu}_{i}^{c}=e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}(\frac{(\tilde{\nu}_{i}^{c})^{\Re}+i(\tilde{\nu}_{i}^{c})^{\Im}}{\sqrt{2}}+\upsilon_{\nu_{i}^{c}}),$ (4)
and
$\tan\beta={\upsilon_{u}\over{\sqrt{\upsilon_{d}^{2}+\upsilon_{\nu_{i}}\upsilon_{\nu_{i}}}}}\approx{\upsilon_{u}\over\upsilon_{d}},\qquad\upsilon=\sqrt{\upsilon_{u}^{2}+\upsilon_{d}^{2}+\upsilon_{\nu_{i}}\upsilon_{\nu_{i}}}\approx\sqrt{\upsilon_{u}^{2}+\upsilon_{d}^{2}},$ (5)
In supersymmetric (SUSY) extensions of the SM, the R-parity of a particle is defined as $R=(-1)^{L+3B+2S}$ MSSM ; MSSM1 ; MSSM2 ; MSSM3 ; MSSM4 . R-parity is violated if either the baryon number ($B$) or lepton number ($L$) is not conserved, where $S$ denotes the spin of concerned component field. The last two terms in Eq. (1) explicitly violate lepton number and R-parity. For example, if one assigns $L=1$ to the right-handed neutrino superfields, then the last term ${1\over 3}\kappa_{ijk}\hat{\nu}_{i}^{c}\hat{\nu}_{j}^{c}\hat{\nu}_{k}^{c}$ of Eq. (2) violates the lepton number by three units contrary to the $\hat{L}_{i}\hat{L}_{j}\hat{e}^{c}_{k}$ term of the R-parity violating MSSM which shows $\Delta L=1$ effect. R-parity breaking implies that the lightest supersymmetric particle (LSP) is no longer stable.
II.1 The $\mu\nu$SSM Higgs potential
The neutral scalar potential of the tree-level can be written as:
$V^{0}=V_{soft}+V_{D}+V_{F}$ (6)
with
$V_{soft}\!\!\!\!\!$ $=m_{H_{d}}^{2}H_{d}^{0}H_{d}^{0\ast}+m_{H_{u}}^{2}H_{u}^{0}H_{u}^{0\ast}+m_{\tilde{L}_{ij}}^{2}\tilde{\nu}_{i}\tilde{\nu}_{j}^{*}+m_{\tilde{\nu}_{ij}^{c}}^{2}$ (7)
$-((A_{\lambda}e^{i\phi_{\lambda}}\lambda)_{i}\tilde{\nu}_{i}^{c}H_{d}^{0}H_{u}^{0*}-{{1}\over{3}}(A_{\kappa}\kappa)_{ijk}\tilde{\nu}_{i}^{c}\tilde{\nu}_{j}^{c}\tilde{\nu}_{k}^{c}+h.c.)$
$V_{D}={{G^{2}}\over{8}}(\tilde{\nu}_{i}\tilde{\nu}_{i}^{*}+H_{d}^{0}H_{d}^{0*}-H_{u}^{0}H_{u}^{0*})^{2}$ (8)
$V_{F}\!\!\!\!\!$ $=\lambda_{i}\lambda_{i}^{*}H_{d}^{0}H_{d}^{0*}H_{u}^{0}H_{u}^{0*}+e^{i\phi_{\lambda_{i}}}e^{-i\phi_{\lambda_{j}}}\lambda_{i}\lambda_{j}^{*}e^{-i\phi_{\tilde{\nu}_{i}^{c}}}e^{-i\phi_{\tilde{\nu}_{j}^{c}}}\tilde{\nu}_{i}^{c*}\tilde{\nu}_{j}^{c*}(H_{d}^{0}H_{d}^{0*}+H_{u}^{0}H_{u}^{0*})$ (9)
$+\kappa_{ijk}\kappa_{ljm}^{*}e^{i\phi_{\tilde{\nu}_{i}^{c}}}e^{i\phi_{\tilde{\nu}_{k}^{c}}}e^{-i\phi_{\tilde{\nu}_{l}^{c}}}e^{-i\phi_{\tilde{\nu}_{m}^{c}}}\tilde{\nu}_{i}^{c}\tilde{\nu}_{k}^{c}\tilde{\nu}_{l}^{c*}\tilde{\nu}_{m}^{c*}-(e^{-i\phi_{\lambda_{j}}}\lambda_{j}^{*}\kappa_{ijk}\nu_{i}^{c}\nu_{k}^{c}H_{d}^{0}H_{u}^{0*}+h.c.)$
where $G^{2}=g_{1}^{2}+g_{2}^{2}$. The tree-level neutral scalar potential include the usual soft terms and $D-$ and $F-$term contributions. In this work, we take all parameters in the potential ara real.
By using the effective potential methods epm ; epm1 ; epm2 ; epm3 ; epm4 ; epm5 ; epm6 ; epm7 ; epm8 ; epm9 ; epm10 ; epm11 ; epm12 ; epm13 ; epm14 ; epm15 ; epm16 , one can get the one-loop efffective potential:
$V^{1}={{1}\over{32\pi^{2}}}\left\{\sum_{\tilde{f}}N_{f}m_{\tilde{f}}^{4}\left(\log{{m_{\tilde{f}}^{2}}\over{Q^{2}}}-{{3}\over{2}}\right)-2\sum_{f=t,b,\tau}N_{f}m_{f}^{4}\left(\log{{m_{f}^{2}}\over{Q^{2}}}-{{3}\over{2}}\right)\right\}.$ (10)
Here, $Q$ is the renormalization scale, $N_{t}=N_{b}=3$ and $N_{\tau}=1$. $f=t,b,\tau$ denote the third fermions and $\tilde{f}=\tilde{t}_{1,2}=\tilde{b}_{1,2}=\tilde{\tau}_{1,2}$ are the corresponding supersymmetric partners.
Considering the one-loop effective potential, the Higgs potential can be writtrn as:
$V=V^{0}+V^{1}$ (11)
we can calculte the minimization conditions of the potential and the Higgs masses in this work.
II.2 Higgs masses and CPV phases
In the $\mu\nu$SSM, the left- & right-handed sneutrino VEVs lead to the mixing of the neutral components of the Higgs doublets with the left- &\&& right-handed sneutrinos producing an $8\times 8$ CP-even neutral scalar mass matrix, which can be seen in Refs. mnSSM1 ; mnSSM1-1 ; Zhang1 ; MASS . The mixing gives a rich phenomenology in the Higgs sector of the $\mu\nu$SSM mnSSM1 ; mnSSM2 ; mnSSM1-1 ; parameter-space ; mnSSM2-1 ; HZrr ; muon ; MASS ; HZr .
Here, we note that the Higgs doublets and right-handed sneutrinos are basically decoupled from the left-handed sneutrinos MASS , so we didn’t considering the left-handed sneutrinos in the CP-even and CP-odd scalars part.
The CP-even sector mix with the CP-odd sector, the 10 $\times$ 10 mixing matrix is defined by
$M_{h}^{2}=\left(\begin{array}[]{cc}M_{S}^{2}&M_{SP}^{2}\\ (M_{SP}^{2})^{2}&M_{P}^{2}\\ \end{array}\right)$ (14)
with $M_{S}^{2}$ denote the CP-even neutral scalars, $M_{P}^{2}$ is the CP-odd neutral scalars, $M_{SP}^{2}$ represents the mass submatrix for the mixing of CP-even neutral scalars and CP-odd neutral scalars.
The mass sqared matrix $M_{h}^{2}$ can be diagonalized as:
$Z_{H}M_{h}^{2}Z_{H}^{T}=m_{h}^{2}$ (15)
with CPV in the CP-even and CP-odd scalar sector the matrix $Z_{H}$ can be complex.
We consider the radiative correcttions in mass submatrix $M_{H}^{2}$, the radiative corrections from the third fermions ($f=t,b,\tau$) and their superpartners include the two-loop leading-log effects leading-log ; leading-log1 ; leading-log2 . The CP-even neutral scalars is given as:
$M_{S}^{2}=\left(\begin{array}[]{cc}M_{H}^{2}&M_{Mix}^{2}\\ (M_{Mix}^{2})^{2}&M_{\tilde{\nu}_{R}^{c}}^{2}\\ \end{array}\right)$ (18)
In detail, the mass submatrix $M_{H}^{2}$ is defind by
$M_{H}^{2}=\left(\begin{array}[]{cc}M_{h_{d}h_{d}}^{2}+\Delta_{11}&M_{h_{d}h_{u}}^{2}+\Delta_{12}\\ M_{h_{d}h_{u}}^{2}+\Delta_{12}&M_{h_{u}h_{u}}^{2}+\Delta_{22}\\ \end{array}\right)$ (21)
The dominating contributions of radiative corrections $\Delta_{11},\Delta_{12}$ and $\Delta_{22}$ comes from the fermions ($t,b$) and their superpartners
$\Delta_{11}=\Delta_{11}^{t}+\Delta_{11}^{b}$
$\Delta_{12}=\Delta_{12}^{t}+\Delta_{12}^{b}$
$\Delta_{22}=\Delta_{22}^{t}+\Delta_{22}^{b}$ (22)
We didn’t consider the terms containing coupling $Y_{\nu_{i}}$ and $\upsilon_{\nu_{i}}$, because these terms are very small. The radiative corrections from the top quark is give by HZrr ; radiative corr ; radiative corr1 ; radiative corr2 ; radiative corr3 ; radiative corr4 ; radiative corr5
$\Delta_{11}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}{\beta}}}{{(\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu-|\mu|^{2}\cot\beta)^{2}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (23)
$\Delta_{12}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}{\beta}}}{{|\mu|^{2}\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (24)
×$\times\left(\ln{{m_{\tilde{t}_{1}^{2}}^{2}}\over{m_{\tilde{t}_{2}}^{2}}}+{{\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\right)$
$\Delta_{22}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}{\beta}}}\Bigg{\{}\ln\frac{m_{\tilde{t}_{1}}^{2}m_{\tilde{t}_{2}}^{2}}{m_{t^{4}}}+{{2|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cos\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}\ln{{m_{\tilde{t}_{1}^{2}}^{2}}\over{m_{\tilde{t}_{2}}^{2}}}$ (25)
$+{{(2|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta)^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2})^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})+{{1}\over{16\pi^{2}}}\ln{{m_{\tilde{t}_{1}}^{2}m_{\tilde{t}_{2}}^{2}}\over{m_{t}^{4}}}$
×$\times\left({{3e^{2}m_{t}^{2}}\over{4s_{w}^{2}m_{W}^{2}}}-32\pi\alpha_{s}\right)\Bigg{[}{1\over 2}\ln{{m_{\tilde{t}_{1}}^{2}m_{\tilde{t}_{2}}^{2}}\over{m_{t}^{4}}}+{{2|A_{t}-\mu\cot\beta|^{2}}\over{m_{\tilde{t}_{1}}m_{\tilde{t}_{2}}}}\Bigg{(}1-{{(A_{t}-\mu\cot\beta)^{2}}\over{12m_{\tilde{t}_{1}}^{2}m_{\tilde{t}_{2}}^{2}}}\Bigg{)}\Bigg{]}\Bigg{\}}$
with
$g(m_{1}^{2},m_{2}^{2})=2-{{m_{1}^{2}+m_{2}^{2}}\over{m_{1}^{2}-m_{2}^{2}}}\ln{{m_{1}^{2}}\over{m_{2}^{2}}}$ (26)
to save space, the mass matrix and the radiative corrections are given in the Appendix A.
In the radiative corrections, the trilinear coupling $A_{t}=|A_{t}|e^{i\phi_{A_{t}}}$ can be complex. These seven independent phase which have been defined as:
$\phi_{\lambda},\phi_{A_{t}},\phi_{\upsilon_{u}},\phi_{\upsilon_{d}},\phi_{\upsilon_{\nu_{i}^{c}}}$ (27)
III Excess at 95 GeV
[…]
$\mu_{LEP}^{bb}={{\sigma(e^{+}e^{-}\rightarrow Zh_{1}\rightarrow Zb\bar{b})}\over{\sigma^{SM}(e^{+}e^{-}\rightarrow Zh_{1}\rightarrow Zb\bar{b})}}=0.117\pm 0.057$ (28)
The value for $\mu_{LEP}^{bb}$ can be found in 98LEP1 ; 98LHC with the method introduced in 98LEP2 . $h_{1}$ is the Higgs which mass range around $\sim$ 96 GeV and $h_{2}$ is the 125 GeV Higgs boson in the following. In the feamework of the $\mu\nu$SSM we use $\mu_{NP}^{bb}$ to describe the signal strength, the expression for $\mu_{NP}^{bb}$ can be approximated as 98LEP3
$\mu_{NP}^{bb}\!\!\!\!\!$ $={{\sigma^{\text{NP}}(Z^{*}\rightarrow Zh_{1})}\over{\sigma^{\text{SM}}(Z^{*}\rightarrow Zh_{1})}}\times{{Br^{\text{NP}}(h_{1}\rightarrow b\bar{b})}\over{Br^{\text{SM}}(h_{1}\rightarrow b\bar{b})}}$ (29)
$\approx|C_{h_{1}VV}|^{2}\times{{\Gamma^{\text{NP}}_{h_{1}\rightarrow b\bar{b}}}\over{\Gamma_{h_{1}\rightarrow b\bar{b}}^{\text{SM}}}}\times{{\Gamma_{\text{tot}}^{\text{SM}}}\over{\Gamma_{\text{tot}}^{\text{NP}}}}$
$\approx{{|C_{h_{1}VV}|^{2}\times|C_{h_{1}d\bar{d}}|^{2}}\over{|C_{h_{1}d\bar{d}}|^{2}(Br_{h_{1}\rightarrow b\bar{b}}^{\text{SM}}+Br_{h_{1}\rightarrow\tau\bar{\tau}}^{\text{SM}})+|C_{h_{1}u\bar{u}}|^{2}(Br_{h_{1}\rightarrow g\bar{g}}^{\text{SM}}+Br_{h_{1}\rightarrow c\bar{c}}^{\text{SM}})}}$
One can found the SM branching ratios $Br^{SM}$ in Ref. PDG1 , $\Gamma$ is the decay widths. The couplings are normalized to the SM-prediction of a Higgs boson of the same mass. $C_{h_{1}}$ is the coupling of $h_{1}$ and gauge boson, $C_{h_{1}u\bar{u}}$ and $C_{h_{1}d\bar{d}}$ are the couplings of $h_{1}$ and up- and down-type quarks. The normalized couplings are given in:
$C_{h_{i}d\bar{d}}={{Z_{H}^{i1}}\over{\cos\beta}},\qquad C_{h_{i}u\bar{u}}={{Z_{H}^{i2}}\over{\sin\beta}},\qquad C_{h_{i}VV}=Z_{H}^{i1}\cos\beta+Z_{H}^{i2}\sin\beta,$ (30)
[…]
$\mu_{\gamma\gamma}^{CMS}={{\sigma(gg\rightarrow h_{1}\rightarrow\gamma\gamma)}\over{\sigma^{SM}(gg\rightarrow h_{1}\rightarrow\gamma\gamma)}}=0.6\pm 0.2,$ (31)
Recently, ATLAS reported their new results at 95.4 GeV based on full Run 2 data set 95ATLAS , the “di-photon excess” with a signal strength of
$\mu_{\gamma\gamma}^{ATLAS}=0.18\pm 0.1$ (32)
Meanwhile, the corresponding CMS result for the “di-photon excess” is given by CMS4
$\mu_{\gamma\gamma}^{CMS}={{\sigma^{exp}(pp\rightarrow\phi\rightarrow\gamma\gamma)}\over{\sigma^{SM}(pp\rightarrow H_{SM}\rightarrow\gamma\gamma)}}=0.33_{-0.12}^{+0.19}.$ (33)
neglecting possible correlations one can get a combined signal strength of CMS5
$\mu_{\gamma\gamma}^{exp}=\mu_{\gamma\gamma}^{ATLAS+CMS}=0.24_{-0.08}^{+0.09}.$ (34)
In this work, the approximation of the diphoton rate of the $h_{1}$ can written as 98LEP1 ; 98LEP3 :
$\mu_{NP}^{\gamma\gamma}\!\!\!\!\!$ $={{\sigma^{\text{NP}}(gg\rightarrow h_{1})}\over{\sigma^{\text{SM}}(gg\rightarrow h_{1})}}\times{{Br^{\text{NP}}(h_{1}\rightarrow\gamma\gamma)}\over{Br^{\text{SM}}(h_{1}\rightarrow\gamma\gamma)}}$ (35)
$\approx|C_{h_{1}u\bar{u}}|^{2}\times{{\Gamma_{h_{1}\rightarrow\gamma\gamma}^{\text{NP}}}\over{\Gamma_{h_{1}\rightarrow\gamma\gamma}^{\text{SM}}}}\times{{\Gamma_{\text{tot}}^{\text{SM}}}\over{\Gamma_{\text{tot}}^{\text{NP}}}}$
$\approx{{|C_{h_{1}u\bar{u}}|^{2}\times|C_{h_{1}\gamma\gamma}|^{2}}\over{|C_{h_{1}d\bar{d}}|^{2}(Br_{h_{1}\rightarrow b\bar{b}}^{\text{SM}}+Br_{h_{1}\rightarrow\tau\bar{\tau}}^{\text{SM}})+|C_{h_{1}u\bar{u}}|^{2}(Br_{h_{1}\rightarrow gg}^{\text{SM}}+Br_{h_{1}\rightarrow c\bar{c}}^{\text{SM}})}}$
The effective coupling $C_{h_{1}\gamma\gamma}$ can be written as 98LEP3
$|C_{h_{1}\gamma\gamma}|^{2}={{|{4\over 3}C_{h_{1}t\bar{t}}A_{1/2}(\tau_{t})+C_{h_{1}VV}A_{1}(\tau_{W})|^{2}}\over{|{4\over 3}A_{1/2}(\tau_{t})+A_{1}(\tau_{W})|^{2}}}.$ (36)
with $\tau_{t}={{m_{h_{1}}^{2}}\over{4m_{t}^{2}}}<1$ and $\tau_{t}={{m_{h_{1}}^{2}}\over{4m_{W}^{2}}}<1$, the from factor $A_{1/2}$ and $A_{1}$ are given by Chrr
$A_{1/2}(x)\!\!\!\!\!$ $=2(x+(x-1)\arcsin^{2}\sqrt{x})x^{-2},\qquad\qquad\qquad x\leq 1$ (37)
$A_{1}(x)\!\!\!\!\!$ $=-(2x^{2}+3x+3(2x-1)\arcsin^{2}\sqrt{x})x^{-2}\qquad x\leq 1$ (38)
By using Eq.(29) and Eq.(35), we calculate the two signal strengths.
IV Numerical Results
In this section we will discuss the couplings and signal strength of 96 GeV Higgs in CPV $\mu\nu$SSM. The free parameters in our analysis will be
$\lambda,\tan\beta,\kappa,A_{t},A_{\lambda},A_{\kappa},A_{b},\upsilon_{\nu^{c}}$ (39)
we take $\upsilon_{\nu_{1}^{c}}=\upsilon_{\nu_{2}^{c}}=\upsilon_{\nu_{3}^{c}}$, and we have defined :
${\kappa_{ijk}}=\kappa{\delta_{ij}}{\delta_{jk}},\quad{({A_{\kappa}}\kappa)_{ijk}}={A_{\kappa}}\kappa{\delta_{ij}}{\delta_{jk}},\quad\lambda_{i}=\lambda,$
${({A_{\lambda}}\lambda)}_{i}={A_{\lambda}}\lambda,\quad{Y_{{e_{ij}}}}={Y_{{e_{i}}}}{\delta_{ij}},\quad{({A_{e}}{Y_{e}})_{ij}}={A_{e}}{Y_{{e_{i}}}}{\delta_{ij}},$
${Y_{{\nu_{ij}}}}={Y_{{\nu_{i}}}}{\delta_{ij}},\quad(A_{\nu}Y_{\nu})_{ij}={a_{{\nu_{i}}}}{\delta_{ij}},\quad m_{\tilde{\nu}_{ij}^{c}}^{2}=m_{\tilde{\nu}_{i}^{c}}^{2}{\delta_{ij}},$
$m_{\tilde{Q}_{ij}}^{2}=m_{{{\tilde{Q}_{i}}}}^{2}{\delta_{ij}},\quad m_{\tilde{u}_{ij}^{c}}^{2}=m_{{{\tilde{u}_{i}}^{c}}}^{2}{\delta_{ij}},\quad m_{\tilde{d}_{ij}^{c}}^{2}=m_{{{\tilde{d}_{i}}^{c}}}^{2}{\delta_{ij}},$
$m_{{{\tilde{L}}_{ij}}}^{2}=m_{{\tilde{L}}}^{2}{\delta_{ij}},\quad m_{\tilde{e}_{ij}^{c}}^{2}=m_{{{\tilde{e}}^{c}}}^{2}{\delta_{ij}},\quad\upsilon_{\nu_{i}^{c}}=\upsilon_{\nu^{c}},$ (40)
where $i,j=1,2,3$
IV.1 Mass and Coupling
Figure 1: Values for $m_{h_{2}}$ versus $m_{h_{2}}$ in (a), the normalized couplings $|C_{h_{2}VV}|$ versus $|C_{h_{1}VV}|$ in (b). The input parameters in Tab. 1.
[…]
Figure 2: Values for $m_{h_{1}}$ versus the parameters $\upsilon_{\nu^{c}}$ in (a) and $\kappa$ in (c), values for $m_{h_{2}}$ versus the parameters $\upsilon_{\nu^{c}}$ in (b) and $\kappa$ in (d). We take $\tan\beta=5$ and other parameters in Tab. 1.
In Fig.2, the left column, we can see that $m_{h_{1}}$ is very sensitive to $\upsilon_{\nu^{c}}$. As $\upsilon_{\nu^{c}}$ becomes larger, $m_{h_{1}}$ will rapidly become smaller, on the contrary, $m_{h_{2}}$ will slowly become larger. In order for $m_{h_{1}}$ to be around 95.4 GeV, the value of $\upsilon_{\nu^{c}}$ will not have a wide range. Similar to $\upsilon_{\nu^{c}}$, $m_{h_{1}}$ is also very sensitive to $\kappa$, and with the increase of $\lambda$, a larger $\kappa$ value can be taken to keep $m_{h_{1}}$ near 95.4 GeV. On the contrary, $\kappa$ has little effect on $m_{h_{2}}$, especially when $\lambda$=0.085 or 0.087, $m_{h_{2}}$ will be very stable.
$\qquad A_{\lambda}\qquad$ $\qquad\lambda\qquad$ $\qquad A_{t}\qquad$ $\qquad\upsilon_{\nu^{c}}\qquad$ $\qquad\kappa\qquad$ $\qquad A_{\kappa}\qquad$
920 [0.08;0.09] 1660 1000 0.042 -370
$\qquad A_{b}\qquad$ $\qquad\phi_{\lambda}/\pi\qquad$ $\qquad\phi_{A_{t}}/\pi\qquad$ $\qquad\phi_{v_{u}}/\pi\qquad$ $\qquad\phi_{v_{d}}/\pi\qquad$ $\qquad\phi_{\upsilon_{\nu^{c}}}/\pi\qquad$
700 -0.020 -0.200 0.380 -0.095 -1
Table 1: Input parameters to fit the LEP and the CMS excesses. All dimensionful parameters are given in GeV.
Figure 3: The component of $h_{1}$ (a) and $h_{2}$ (b). The input parameters in Tab.1, and we take $\tan\beta=4.31$.
In Fig. 3, we have showed the component of $h_{1}$ and $h_{2}$, for $h_{1}$, CP-odd $H_{u}$ component and CP-even right-handed sneutrinos component are the main components, and with the increase of $\lambda$, CP-odd $H_{d}$ component will become smaller, while the CP-even right-handed sneutrino component will become larger. While for $h_{2}$, CP-even $H_{u}$ is the main component, as $\lambda$ increases, the CP-even $H_{u}$ component will increase, the CP-even $H_{d}$ component will gradually decrease, and the CP-odd $H_{d}$ component will gradually increase.
Figure 4: The left column is the correlation of $h_{1}$ and the CPV phases, the grey area represents the experimental error of 3 GeV. The right column is the correlation of $h_{2}$ and the CPV phases, the grey area is the theory uncertainty of 1 GeV.
Then, in Fig. 4, we analyze the correlation between the Higgs masses and the CPV phases. For the first row, we take $\lambda=0.09,\tan\beta=6,\kappa=0.315,$ for the second row, $\lambda=0.095,\tan\beta=4.5,\kappa=0.315,$ for the third row, $\lambda=0.098,\tan\beta=4.3,\kappa=0.32$, for the last row, $\lambda=0.085,\tan\beta=3,\kappa=0.04$. Other parameter keep same as Tab.1. We should remark that there are many CPV phase values that can be constrained by the Higgs masses of Fig.7 (a), and here we only select one of them. The signal strengths and the Higgs masses do not increase or decrease all the time, but oscillates like a sine function. In Fig. 4, the first row, when $\phi_{\lambda}$ grown from 4 to 5, the mass range of the “di-photon excess” can fit the experimental constraints, the SM-like Higgs mass can be kept around 125 GeV.
[…]
IV.2 Signal Strengths
Figure 5: The left column shows that the signal strength $\mu_{NP}^{bb}$ varies with $\upsilon_{\nu^{c}}$ (a) and $\kappa$ (c), the right column shows that the signal strength $\mu_{NP}^{\gamma\gamma}$ varies with $\upsilon_{\nu^{c}}$ (b) and $\kappa$ (d). We take $\lambda=0.09$ and other parameters in Tab. 1.
[…]
The effects of CPV phases is also very obvious. Here, we take $\phi_{A_{t}}$ as an example, we take $\phi_{A_{t}}$ from $-\pi$ to $\pi$. In the first row of Fig. 6, we can see that the minimum value of $h_{1}$ appears to be around 96.2 GeV and varies periodically with $\phi_{A_{t}}$. The maximum value of $h_{1}$ does not exceed 1 GeV larger than the minimum value. Meanwhile, the highest point of $h_{2}$ is at 125.3 GeV, with a maximum value 1.5GeV larger than the minimum value.
The impact of $\phi_{A_{t}}$ on the normalized couplings are relatively small. In the second row of Fig. 6, we can see that the peak of $|C_{h_{1}VV}|$ is close to 0.313, and the highest point of $|C_{h_{2}VV}|$ is slightly less than 0.856. This implies that if the ”95.4 excess” is a new particle, then its coupling with gauge bosons should be much smaller than the coupling of the SM-like Higgs with gauge bosons.
Figure 6: Higgs masses, couplings, and signal strengths are sensitive to CPV phases. Here, we take $\tan\beta$=4.5, $\lambda$=0.086, $\phi_{A_{t}}=[-\pi,\pi]$, other parameters are the same as Tab. 1.
[…]
$\qquad\lambda\qquad$ $\qquad\tan\beta\qquad$ $\qquad\kappa\qquad$ $\qquad\upsilon_{\nu^{c}}\qquad$ $\qquad A_{\lambda}\qquad$ $\qquad A_{t}\qquad$ $\qquad A_{\kappa}\qquad$
$[0.087;0.090]$ [2.4;4.5] [0.040;0.045] [950;1050] [915;920] [1600;1700] [-375;-370]
$\qquad A_{b}\qquad$ $\qquad\phi_{\lambda}/\pi\qquad$ $\qquad\phi_{A_{t}}/\pi\qquad$ $\qquad\phi_{v_{u}}/\pi\qquad$ $\qquad\phi_{v_{d}}/\pi\qquad$ $\qquad\phi_{\upsilon_{\nu^{c}}}/\pi\qquad$ $\qquad M_{2}\qquad$
$[695;710]$ [-0.024;-0.022] [-0.4;-0.1] [0.382;0.385] [-0.095;-0.090] [-1;-0.9] 800
Table 2: The parameter space of the random scan plot. All dimensionful parameters are given in GeV.
[…]
Let us remark that in the random scan plots, we must first ensure that the mass of $h_{2}$ can conform to the experimental constraints, because the mass determination for SM-like Higgs is already very accurate. In Fig.7 (a), we chose 1 GeV theory uncertainty for the SM-like Higgs. For $h_{1}$, we can choose the theory uncertainty up to 3 GeV, because CMS only shows that there is a 2.8 $\sigma$ excess at 95.3 GeV CMS1 ; CMS5 ; CMS9 , which is not exactly observed. In Fig 7 (b), the $(\mu_{NP}^{\gamma\gamma},\mu_{NP}^{bb})$ plane, we can clearly see that with the parameter space of Tab.2, most points can explain the di-photon excess and the $b\bar{b}$ excess.
V Conclusion
In this paper we introduced CPV in the $\mu\nu$SSM, which lead CP-even Higgs sector mixed with CP-odd Higgs sector. We also analyzed a excess in the di-photon decay mode at $\sim$ 95 GeV as reported by ATLAS and CMS, together with a $\sim$ 2 $\sigma$ excess at LEP in the same mass range. The mixing and CPV are used to produce the lightest Higgs boson mass around 95 GeV, and the next-to-lightest Higgs boson mass around 125 GeV is the so-called SM-like Higgs. The lightest Higgs boson can explain an excess of $\gamma\gamma$ events at $\sim$ 96 GeV as reported by CMS.
In the numerical part, we find a suitable parameter spaces, based on which we show the properties of the lightest and next-to-lightest Higgs boson. We found it very easy get both $\mu_{NP}^{bb}$ and $\mu_{NP}^{\gamma\gamma}$ to reach the central values of their respective experiments at the same time. We also analyze the influence of relevant parameters and CPV phases on the signal strengths or Higgs masses, and improve the signal strengths as much as possible while ensuring that the SM-like Higgs meets the experimental constraints and the “di-photon excess” is around 95 GeV.
Acknowledgements.
The work has been supported by the National Natural Science Foundation of China (NNSFC) with Grants No. 12075074, No. 12235008, No. 11705045, No. 11535002, Hebei Natural Science Foundation with Grant No. A2022201017, No. A2023201041, Natural Science Foundation of Guangxi Autonomous Region with Grant No. 2022GXNSFDA035068, the youth top-notch talent support program of the Hebei Province.
Appendix A Msss matyices
The CP-even neutral scalars have the composition $S^{T}=(h_{d},h_{u},(\tilde{\nu}_{i}^{c})^{\Re})$, one can write the mass matrix $M_{S}^{2}$ as:
$M_{S}^{2}=\left(\begin{array}[]{ccccc}M_{h_{d}h_{d}}^{2}&M_{h_{d}h_{u}}^{2}&M_{h_{d}(\tilde{\nu}_{1}^{c})^{\Re}}^{2}&M_{h_{d}(\tilde{\nu}_{2}^{c})^{\Re}}^{2}&M_{h_{d}(\tilde{\nu}_{3}^{c})^{\Re}}^{2}\\ M_{h_{u}h_{d}}^{2}&M_{h_{u}h_{u}}^{2}&M_{h_{u}(\tilde{\nu}_{1}^{c})^{\Re}}^{2}&M_{h_{u}(\tilde{\nu}_{2}^{c})^{\Re}}^{2}&M_{h_{u}(\tilde{\nu}_{3}^{c})^{\Re}}^{2}\\ M_{(\tilde{\nu}_{1}^{c})^{\Re}h_{d}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}h_{u}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Re}}^{2}\\ M_{(\tilde{\nu}_{2}^{c})^{\Re}h_{d}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}h_{u}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Re}}^{2}\\ M_{(\tilde{\nu}_{3}^{c})^{\Re}h_{d}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}h_{u}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Re}}^{2}\\ \end{array}\right)$ (46)
$M_{h_{d}h_{d}}^{2}\!\!\!\!\!$ $={{G^{2}}\over{2}}v_{d}^{2}+\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\tan\beta$ (47)
$+\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\tan\beta+\Delta_{11},$
$M_{h_{d}h_{u}}^{2}\!\!\!\!\!$ $=-{{G^{2}}\over{2}}v_{d}v_{u}-\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}+2|\lambda_{i}|^{2}v_{d}v_{u}$ (48)
$-\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}+\Delta_{12},$
$M_{h_{u}h_{u}}^{2}\!\!\!\!\!$ $={{G^{2}}\over{2}}v_{u}^{2}+\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\cot\beta$ (49)
$+\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\cot\beta+\Delta_{22},$
$M_{h_{d}(\tilde{\nu}_{i}^{c})^{\Re}}^{2}\!\!\!\!\!$ $=-\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}+2\lambda_{i}\lambda_{j}^{*}v_{d}\upsilon_{\nu_{j}^{c}}$ (50)
$-2\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{u}\upsilon_{\nu_{j}^{c}}+\Delta_{1(2+i)},$
$M_{h_{u}(\tilde{\nu}_{i}^{c})^{\Re}}^{2}\!\!\!\!\!$ $=-\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}+2\lambda_{i}\lambda_{j}^{*}v_{u}\upsilon_{\nu_{j}^{c}}$ (51)
$-2\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}\upsilon_{\nu_{j}^{c}}+\Delta_{2(2+i)},$
$M_{(\tilde{\nu}_{i}^{c})^{\Re}(\tilde{\nu}_{i}^{c})^{\Re}}^{2}\!\!\!\!\!$ $=(A_{\kappa}\kappa)_{ijk}\Re(e^{i\phi_{\upsilon_{\nu_{k}^{c}}}}e^{i\phi_{\upsilon_{\nu_{k}^{c}}}}e^{i\phi_{\upsilon_{\nu_{k}^{c}}}})\upsilon_{\nu_{k}^{c}}+\Delta_{(2+i)(2+j)}$ (52)
$+\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}{{v_{d}v_{u}}\over{\upsilon_{\nu_{j}^{c}}}},$
$M_{(\tilde{\nu}_{i}^{c})^{\Re}(\tilde{\nu}_{j}^{c})^{\Re}}^{2}\!\!\!\!\!$ $=|\lambda_{i}|^{2}(v_{d}^{2}+v_{u}^{2})+\Delta_{(2+i)(2+j)}.$ (53)
The rediative corrections from the top quark and bottom quark and their corresponding supersymmetric partners, the corrections in the mass matrix can be expressed as:
$\Delta_{ab}=\Delta_{ab}^{t}+\Delta_{ab}^{b}$ (54)
$\Delta_{1(2+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}\Bigg{\{}{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (55)
×$\times{{|\mu|^{2}\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
$\Delta_{2(2+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}\Bigg{\{}{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}\Bigg{\}}$ (56)
×$\times\Bigg{\{}\ln{{m_{\tilde{t}_{1}}^{2}}\over{m_{\tilde{t}_{2}}^{2}}}+{{|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
$\Delta_{(2+i)(2+j)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}\Bigg{\{}{{({1\over 2}\upsilon_{\nu_{j}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d})^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2})^{2}}}$ (57)
×$\times g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}}$
$\Delta_{11}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}\ln{{m_{\tilde{b}_{1}}^{2}m_{\tilde{b}_{2}}^{2}}\over{m_{b}^{4}}}+{{2|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\ln{{m_{\tilde{b}_{1}}^{2}}\over{m_{\tilde{b}_{2}}^{2}}}$ (58)
$+{{(|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
$\Delta_{12}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$ (59)
×$\times\Bigg{\{}\ln{{m_{\tilde{b}_{1}}^{2}}\over{m_{\tilde{b}_{2}}^{2}}}+{{|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
$\Delta_{22}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{(|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),\Bigg{\}}$ (60)
$\Delta_{1(2+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$ (61)
×$\times\Bigg{\{}\ln{{m_{\tilde{b}_{1}}^{2}}\over{m_{\tilde{b}_{2}}^{2}}}+{{|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
$\Delta_{2(2+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (62)
×$\times{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$
$\Delta_{(2+i)(2+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{({1\over 2}\upsilon_{\nu_{j}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u})^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}$ (63)
×$\times g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}).$
In the same way, the CP-odd neutral scalars mass matrix is that:
$M_{P}^{2}=\left(\begin{array}[]{ccccc}M_{\sigma_{d}\sigma_{d}}^{2}&M_{\sigma_{d}\sigma_{u}}^{2}&M_{\sigma_{d}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{\sigma_{d}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{\sigma_{d}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{\sigma_{u}\sigma_{d}}^{2}&M_{\sigma_{u}\sigma_{u}}^{2}&M_{\sigma_{u}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{\sigma_{u}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{\sigma_{u}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{1}^{c})^{\Im}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Im}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Im}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Im}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Im}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{2}^{c})^{\Im}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Im}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Im}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Im}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Im}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{3}^{c})^{\Im}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Im}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Im}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Im}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Im}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ \end{array}\right)$ (69)
$M_{\sigma_{d}\sigma_{d}}^{2}\!\!\!\!\!$ $=\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\tan\beta$ (70)
$+\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\tan\beta+\Delta_{66},$
$M_{\sigma_{d}\sigma_{u}}^{2}\!\!\!\!\!$ $=-\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}$ (71)
$+\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}+\Delta_{67},$
$M_{\sigma_{u}\sigma_{u}}^{2}\!\!\!\!\!$ $=\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\cot\beta$ (72)
$+\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\cot\beta+\Delta_{77},$
$M_{\sigma_{d}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}$ (73)
$+2\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{u}\upsilon_{\nu_{i}^{c}}+\Delta_{6(7+i)},$
$M_{\sigma_{u}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=-\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}$ (74)
$-2\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}\upsilon_{\nu_{i}^{c}}+\Delta_{7(7+i)},$
$M_{(\tilde{\nu}_{i}^{c})^{\Im}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}{{v_{d}v_{u}}\over{\upsilon_{\nu_{i}^{c}}}}-(A_{\kappa}\kappa)_{ijk}\Re(e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})\upsilon_{\nu_{k}^{c}}$ (75)
$+4\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}v_{u}+\Delta_{(7+i)(7+i)},$
$M_{(\tilde{\nu}_{i}^{c})^{\Im}(\tilde{\nu}_{j}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=|\lambda_{i}|^{2}(v_{u}^{2}+v_{d}^{2})+\Delta_{(7+i)(7+i)}.$ (76)
$\Delta_{66}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{(-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu)^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}})^{2}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$ (77)
$\Delta_{67}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (78)
×$\times{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
$\Delta_{77}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{(\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta)^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2})^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$ (79)
$\Delta_{6(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (80)
×$\times{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
$\Delta_{7(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (81)
×$\times{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
$\Delta_{(7+i)(7+j)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{(-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d})^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}})^{2}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (82)
$\Delta_{66}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(-\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}})^{2}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (83)
$\Delta_{67}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-(\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu)^{2}\tan\beta}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (84)
$\Delta_{77}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (85)
$\Delta_{6(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-(\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b})^{2}\mu\tan\beta\lambda_{i}v_{u}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (86)
$\Delta_{7(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b})^{2}\mu\lambda_{i}v_{u}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (87)
$\Delta_{(7+i)(7+j)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u})^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}).$ (88)
The mix mass submatrix $M_{SP}^{2}$
$M_{SP}^{2}=\left(\begin{array}[]{ccccc}M_{h_{d}\sigma_{d}}^{2}&M_{h_{d}\sigma_{u}}^{2}&M_{h_{d}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{h_{d}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{h_{d}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{h_{u}\sigma_{d}}^{2}&M_{h_{u}\sigma_{u}}^{2}&M_{h_{u}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{h_{u}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{h_{u}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{1}^{c})^{\Re}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{2}^{c})^{\Re}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{3}^{c})^{\Re}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ \end{array}\right)$ (94)
$M_{h_{d}\sigma_{d}}^{2}\!\!\!\!\!$ $=\Delta_{16},$ (95)
$M_{h_{d}\sigma_{u}}^{2}\!\!\!\!\!$ $=\Delta_{17},$ (96)
$M_{h_{u}\sigma_{d}}^{2}\!\!\!\!\!$ $=\Delta_{26},$ (97)
$M_{h_{u}\sigma_{u}}^{2}\!\!\!\!\!$ $=\Delta_{27},$ (98)
$M_{h_{d}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=-\Im(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}+\Delta_{1(7+i)},$ (99)
$M_{h_{u}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=-\Im(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}+\Delta_{2(7+i)},$ (100)
$M_{(\tilde{\nu}_{i}^{c})^{\Re}\sigma_{d}}^{2}\!\!\!\!\!$ $=-\Im(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}+\Delta_{(2+i)6},$ (101)
$M_{(\tilde{\nu}_{i}^{c})^{\Re}\sigma_{u}}^{2}\!\!\!\!\!$ $=\Im(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}+\Delta_{(2+i)7},$ (102)
$M_{(\tilde{\nu}_{i}^{c})^{\Re}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=2\Im(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}v_{u}$ (103)
$-(A_{\kappa}\kappa)_{i}\Re(e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})\upsilon_{\nu_{i}^{c}}+\Delta_{(2+i)(7+i)},$
$M_{(\tilde{\nu}_{i}^{c})^{\Re}(\tilde{\nu}_{j}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=\Delta_{(2+i)(7+i)}.$ (104)
$\Delta_{16}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{|\mu|^{2}\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (105)
×$\times{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
$\Delta_{17}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{|\mu|^{2}\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (106)
×$\times{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\tan\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
$\Delta_{26}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (107)
×$\times\Bigg{\{}\ln{m_{\tilde{t}_{1}}^{2}\over m_{\tilde{t}_{2}}^{2}}+{{|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
$\Delta_{27}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (108)
×$\times\Bigg{\{}\ln{m_{\tilde{t}_{1}}^{2}\over m_{\tilde{t}_{2}}^{2}}+{{|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
$\Delta_{1(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{|\mu|^{2}\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (109)
×$\times{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
$\Delta_{2(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (110)
×$\times\Bigg{\{}\ln{m_{\tilde{t}_{1}}^{2}\over m_{\tilde{t}_{2}}^{2}}+{{|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
$\Delta_{(2+i)6}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (111)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
$\Delta_{(2+i)7}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (112)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}},$
$\Delta_{(2+i)(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (113)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}},$
$\Delta_{16}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}$ (114)
×$\times\Bigg{\{}\ln{m_{\tilde{b}_{1}}^{2}\over m_{\tilde{b}_{2}}^{2}}+{{|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
$\Delta_{17}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (115)
×$\times\Bigg{\{}\ln{m_{\tilde{b}_{1}}^{2}\over m_{\tilde{b}_{2}}^{2}}+{{|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}},$
$\Delta_{26}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}$ (116)
×$\times{{|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$
$\Delta_{27}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (117)
×$\times{{|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
$\Delta_{1(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (118)
×$\times\Bigg{\{}\ln{m_{\tilde{b}_{1}}^{2}\over m_{\tilde{b}_{2}}^{2}}+{{|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$
$\Delta_{2(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (119)
×$\times{{|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
$\Delta_{(2+i)6}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (120)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
$\Delta_{(2+i)7}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (121)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
$\Delta_{(2+i)(7+j)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (122)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}.$

### 06 Mar 2024 v2

@@
 I Introduction
-In 2012, the 125 GeV Higgs boson discovered by the Large Hadron Collider (LHC) mh-ATLAS ; mh-CMS , the measured mass of the Higgs boson now is PDG1 $m_{h}=125.25\pm 0.17\>{\rm{GeV}}.$ The discovery of the Higgs boson marked a huge success for the Standard Model (SM), but it did not stop the search for New Physics (NP) at the LHC, and one of them was the search for lighter scalar particles.
+In 2012, the 125 GeV Higgs boson discovered by the Large Hadron Collider (LHC) mh-ATLAS ; mh-CMS ; the measured mass of the Higgs boson now is PDG1 $m_{h}=125.25\pm 0.17\>{\rm{GeV}}$.
+The discovery of the Higgs boson marked a huge success for the standard model (SM), but it did not stop the search for new physics (NP) at the LHC, and one of them was the search for lighter scalar particles.
 […]
-As one of the extensions of the SM, the $\mu$ from $\nu$ supersymmetric standard model ($\mu$$\nu$SSM) mnSSM ; mnSSM1 ; mnSSM1-1 ; parameter-space ; mnSSM2 ; mnSSM2-1 ; Zhang1 ; Zhang2 can solve the $\mu$ problem m-problem of the minimal supersymmetric standard model (MSSM) MSSM ; MSSM1 ; MSSM2 ; MSSM3 ; MSSM4 , through introducing three singlet right-handed neutrino superfields $\hat{\nu}_{i}^{c}$ ($i=1,2,3$).
+As one of the extensions of the SM, the $\mu$-from-$\nu$ supersymmetric standard modelmnSSM ; mnSSM1 ; mnSSM1-1 ; parameter-space ; mnSSM2 ; mnSSM2-1 ; Zhang1 ; Zhang2 can solve the $\mu$ problem m-problem of the minimal supersymmetric standard model (MSSM) MSSM ; MSSM1 ; MSSM2 ; MSSM3 ; MSSM4 , through introducing three singlet right-handed neutrino superfields $\hat{\nu}_{i}^{c}$ ($i=1,2,3$).
 The neutrino superfields lead the mixing of the neutral components of the Higgs doublets with the right-handed sneutrinos, that is different from the Higgs sector of the MSSM.
-The mixing can change the Higgs couplings, and influence the decay processes of the Higgs bosons.
-In addition, we also introduce CP-violation, and we also get a lighter Higgs at $\sim$ 95 GeV with a suitable parameter space.
+The mixing can change the Higgs couplings and influence the decay processes of the Higgs bosons.
+In addition, we also introduce $CP$ violation, and we also get a lighter Higgs at $\sim$ 95 GeV with a suitable parameter space.
 The paper is organized as follows.
 In Sec.
-II, we introduce the CPV $\mu\nu$SSM briefly, about the superpotential and the CPV phases.
+II, we introduce the $CP$-violating $\mu$-from-$\nu$ SSM briefly, about the superpotential and the CPV phases.
 In Sec.
-III, we study the excess at 95 GeV in the CPV $\mu\nu$SSM.
-Sec.
-IV and Sec.
-V, we show the numerical analysis and the conclusion.
+III, we study the excess at 95 GeV in the CPV $\mu$-from-$\nu$ SSM.
+In Secs.
+IV and V, we show the numerical analysis and the conclusion,respectively.
 II The Model
-The superpotential of the $\mu\nu$SSM contains Yukawa couplings for neutrinos, two additional types of terms involving the Higgs doublet superfields $\hat{H}_{u}$ and $\hat{H}_{d}$, and the right-handed neutrino superfields $\hat{\nu}_{i}^{c}$mnSSM
+The superpotential of the $\mu$-from-$\nu$ SSM contains Yukawa couplings for neutrinos, two additional types of terms involving the Higgs doublet superfields $\hat{H}_{u}$ and $\hat{H}_{d}$, and the right-handed neutrino superfields $\hat{\nu}_{i}^{c}$ mnSSM :
 $W={\epsilon_{ab}}\left({Y_{{u_{ij}}}}\hat{H}_{u}^{b}\hat{Q}_{i}^{a}\hat{u}_{j}^{c}+{Y_{{d_{ij}}}}\hat{H}_{d}^{a}\hat{Q}_{i}^{b}\hat{d}_{j}^{c}+{Y_{{e_{ij}}}}\hat{H}_{d}^{a}\hat{L}_{i}^{b}\hat{e}_{j}^{c}\right)$
 $\hskip 27.03003pt+{\epsilon_{ab}}{Y_{{\nu_{ij}}}}\hat{H}_{u}^{b}\hat{L}_{i}^{a}\hat{\nu}_{j}^{c}-{\epsilon_{ab}}{\lambda_{i}}e^{i\phi_{\lambda_{i}}}\hat{\nu}_{i}^{c}\hat{H}_{d}^{a}\hat{H}_{u}^{b}+\frac{1}{3}{\kappa_{ijk}}\hat{\nu}_{i}^{c}\hat{\nu}_{j}^{c}\hat{\nu}_{k}^{c},$ (1)
-where $\hat{H}_{u}^{T}=\Big{(}{\hat{H}_{u}^{+},\hat{H}_{u}^{0}}\Big{)}$, $\hat{H}_{d}^{T}=\Big{(}{\hat{H}_{d}^{0},\hat{H}_{d}^{-}}\Big{)}$, $\hat{Q}_{i}^{T}=\Big{(}{{{\hat{u}}_{i}},{{\hat{d}}_{i}}}\Big{)}$, $\hat{L}_{i}^{T}=\Big{(}{{{\hat{\nu}}_{i}},{{\hat{e}}_{i}}}\Big{)}$ (the index $T$ denotes the transposition) represent the MSSM-like doublet Higgs superfields, and $\hat{u}_{i}^{c}$, $\hat{d}_{i}^{c}$, and $\hat{e}_{i}^{c}$ are the singlet up-type quark, down-type quark and charged lepton superfields, respectively.
-In addition, $Y_{u,d,e,\nu}$, $\lambda$, and $\kappa$ are dimensionless matrices, a vector, and a totally symmetric tensor.
+where $\hat{H}_{u}^{T}=\Big{(}{\hat{H}_{u}^{+},\hat{H}_{u}^{0}}\Big{)}$, $\hat{H}_{d}^{T}=\Big{(}{\hat{H}_{d}^{0},\hat{H}_{d}^{-}}\Big{)}$, $\hat{Q}_{i}^{T}=\Big{(}{{{\hat{u}}_{i}},{{\hat{d}}_{i}}}\Big{)}$, and $\hat{L}_{i}^{T}=\Big{(}{{{\hat{\nu}}_{i}},{{\hat{e}}_{i}}}\Big{)}$ (the index $T$ denotes the transposition) represent the MSSM-like doublet Higgs superfields and $\hat{u}_{i}^{c}$, $\hat{d}_{i}^{c}$, and $\hat{e}_{i}^{c}$ are the singlet up-type quark, down-type quark and charged lepton superfields, respectively.
+In addition, $Y_{u,d,e,\nu}$, $\lambda$, and $\kappa$ are dimensionless matrices, a vector, and a totally symmetric tensor, respectively.
 $a,b=1,2$ are SU(2) indices with antisymmetric tensor $\epsilon_{12}=1$, and $i,j,k=1,2,3$ are generation indices.
-The CP is violated by the parameter $\lambda_{i}$, the CP-violating phase is $\phi_{\lambda_{i}}$.
+The $CP$ is violated by the parameter $\lambda_{i}$, and the $CP$-violating phase is $\phi_{\lambda_{i}}$.
 In the superpotential, if the scalar potential is such that nonzero vacuum expectative values (VEVs) of the scalar components ($\tilde{\nu}_{i}^{c}$) of the singlet neutrino superfields $\hat{\nu}_{i}^{c}$ are induced, the effective bilinear terms $\epsilon_{ab}\varepsilon_{i}\hat{H}_{u}^{b}\hat{L}_{i}^{a}$ and $\epsilon_{ab}\mu\hat{H}_{d}^{a}\hat{H}_{u}^{b}$ are generated, with $\varepsilon_{i}=Y_{\nu_{ij}}\left\langle{\tilde{\nu}_{j}^{c}}\right\rangle$ and $\mu={\lambda_{i}}\left\langle{\tilde{\nu}_{i}^{c}}\right\rangle$, once the electroweak symmetry is broken.
-The last term Eq.
+The last term in Eq.
 (1) generates the effective Majorana masses for neutrinos at the electroweak scale.
-Therefore, the $\mu\nu$SSM can generate three tiny neutrino masses at the tree level through TeV scale seesaw mechanism mnSSM ; neutrino-mass ; neu-mass1 ; neu-mass2 ; neu-mass3 ; neu-mass4 ; neu-mass5 ; neu-mass6 .
-It is worth explaining why TeV scale seesaw was chosen.
-Through a seesaw on the scale of the Grand Unified Theory (GUT), one can get the Yukawa couplings of order one for neutrinos.
+Therefore, the $\mu$-from-$\nu$ SSM can generate three tiny neutrino masses at the tree level through TeV-scale seesaw mechanism mnSSM ; neutrino-mass ; neu-mass1 ; neu-mass2 ; neu-mass3 ; neu-mass4 ; neu-mass5 ; neu-mass6 .
+It is worth explaining why the TeV-scale seesaw was chosen.
+Through a seesaw on the scale of the grand unified theory, one can get Yukawa couplings of the order of one for neutrinos.
 But we know that the Yukawa coupling of the electron is on the order of $10^{-6}$, and the Yukawa couplings of neutrinos can also be around on the order of $10^{-6}$ instead of one.
-In the TeV scale seesaw, this is sufficient to reproduce the neutrino mass, if the Yukawa coupling of the neutrino is of the same order as the Yukawa coupling of the electron mnSSM .
-Here it is important to note that the VEVs of the left-handed sneutrinos $\upsilon_{\nu_{i}}$ are generally small.
-We know that the Dirac masses for the neutrinos $m_{D_{i}}=Y_{\upsilon_{\nu_{i}}}\upsilon_{u}\lesssim 10^{-4}$ GeV in the TeV scale seesaw.
+In the TeV-scale seesaw, this is sufficient to reproduce the neutrino mass, if the Yukawa coupling of the neutrino is of the same order as the Yukawa coupling of the electron mnSSM .
+Here, it is important to note that the VEVs of the left-handed sneutrinos $\upsilon_{\nu_{i}}$ are generally small.
+We know that the Dirac masses for the neutrinos $m_{D_{i}}=Y_{\upsilon_{\nu_{i}}}\upsilon_{u}\lesssim 10^{-4}$ GeV in the TeV-scale seesaw.
 So we can get an estimate of the VEVs of the left-handed sneutrinos, $\upsilon_{\nu_{i}}\lesssim m_{D_{i}}\lesssim 10^{-4}$ GeV, which means that $\upsilon_{\nu_{i}}\ll\upsilon_{d},\upsilon_{u}$ mnSSM ; mnSSM1 .
-The general soft SUSY-breaking terms of the $\mu\nu$SSM are given by
+The general soft supersymmetry-breaking terms of the $\mu$-from-$\nu$SSM are given by
 $-\mathcal{L}_{soft}=m_{{{\tilde{Q}}_{ij}}}^{\rm{2}}\tilde{Q}{{}_{i}^{a\ast}}\tilde{Q}_{j}^{a}+m_{\tilde{u}_{ij}^{c}}^{\rm{2}}\tilde{u}{{}_{i}^{c\ast}}\tilde{u}_{j}^{c}+m_{\tilde{d}_{ij}^{c}}^{2}\tilde{d}{{}_{i}^{c\ast}}\tilde{d}_{j}^{c}+m_{{{\tilde{L}}_{ij}}}^{2}\tilde{L}_{i}^{a\ast}\tilde{L}_{j}^{a}$
@@
 One can define the neutral scalars as
-$H_{d}^{0}=e^{i\phi_{\upsilon_{d}}}(\frac{h_{d}+iP_{d}}{\sqrt{2}}+\upsilon_{d}),\qquad\;\tilde{\nu}_{i}=\frac{(\tilde{\nu}_{i})^{\Re}+i(\tilde{\nu}_{i})^{\Im}}{\sqrt{2}}+\upsilon_{\nu_{i}},$
-$H_{u}^{0}=e^{i\phi_{\upsilon_{u}}}(\frac{h_{u}+iP_{u}}{\sqrt{2}}+\upsilon_{u}),\qquad\tilde{\nu}_{i}^{c}=e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}(\frac{(\tilde{\nu}_{i}^{c})^{\Re}+i(\tilde{\nu}_{i}^{c})^{\Im}}{\sqrt{2}}+\upsilon_{\nu_{i}^{c}}),$ (4)
+$H_{d}^{0}=e^{i\phi_{\upsilon_{d}}}(\frac{h_{d}+iP_{d}}{\sqrt{2}}+\upsilon_{d}),\qquad\;\tilde{\nu}_{i}=\frac{(\tilde{\nu}_{i})^{\text{Re}}+i(\tilde{\nu}_{i})^{\text{Im}}}{\sqrt{2}}+\upsilon_{\nu_{i}},$
+$H_{u}^{0}=e^{i\phi_{\upsilon_{u}}}(\frac{h_{u}+iP_{u}}{\sqrt{2}}+\upsilon_{u}),\qquad\tilde{\nu}_{i}^{c}=e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}(\frac{(\tilde{\nu}_{i}^{c})^{\text{Re}}+i(\tilde{\nu}_{i}^{c})^{\text{Im}}}{\sqrt{2}}+\upsilon_{\nu_{i}^{c}}),$ (4)
 and
 $\tan\beta={\upsilon_{u}\over{\sqrt{\upsilon_{d}^{2}+\upsilon_{\nu_{i}}\upsilon_{\nu_{i}}}}}\approx{\upsilon_{u}\over\upsilon_{d}},\qquad\upsilon=\sqrt{\upsilon_{u}^{2}+\upsilon_{d}^{2}+\upsilon_{\nu_{i}}\upsilon_{\nu_{i}}}\approx\sqrt{\upsilon_{u}^{2}+\upsilon_{d}^{2}},$ (5)
-In supersymmetric (SUSY) extensions of the SM, the R-parity of a particle is defined as $R=(-1)^{L+3B+2S}$ MSSM ; MSSM1 ; MSSM2 ; MSSM3 ; MSSM4 .
-R-parity is violated if either the baryon number ($B$) or lepton number ($L$) is not conserved, where $S$ denotes the spin of concerned component field.
+In supersymmetric extensions of the SM, the $R$ parity of a particle is defined as $R=(-1)^{L+3B+2S}$ MSSM ; MSSM1 ; MSSM2 ; MSSM3 ; MSSM4 .
+$R$ parity is violated if either the baryon number ($B$) or lepton number ($L$) is not conserved, where $S$ denotes the spin of concerned component field.
 The last two terms in Eq.
-(1) explicitly violate lepton number and R-parity.
-For example, if one assigns $L=1$ to the right-handed neutrino superfields, then the last term ${1\over 3}\kappa_{ijk}\hat{\nu}_{i}^{c}\hat{\nu}_{j}^{c}\hat{\nu}_{k}^{c}$ of Eq.
-(2) violates the lepton number by three units contrary to the $\hat{L}_{i}\hat{L}_{j}\hat{e}^{c}_{k}$ term of the R-parity violating MSSM which shows $\Delta L=1$ effect.
-R-parity breaking implies that the lightest supersymmetric particle (LSP) is no longer stable.
-II.1 The $\mu\nu$SSM Higgs potential
-The neutral scalar potential of the tree-level can be written as:
+(1) explicitly violate lepton number and $R$ parity.
+For example, if one assigns $L=1$ to the right-handed neutrino superfields, then the last term ${1\over 3}\kappa_{ijk}\hat{\nu}_{i}^{c}\hat{\nu}_{j}^{c}\hat{\nu}_{k}^{c}$ in Eq.
+(2) violates the lepton number by three units contrary to the $\hat{L}_{i}\hat{L}_{j}\hat{e}^{c}_{k}$ term of the $R$ parity violating MSSM which shows the $\Delta L=1$ effect.
+$R$ parity breaking implies that the lightest supersymmetric particle is no longer stable.
+II.1 The $\mu$-from-$\nu$SSM Higgs potential
+The neutral scalar potential of the tree level can be written as
 $V^{0}=V_{soft}+V_{D}+V_{F}$ (6)
@@
 The tree-level neutral scalar potential include the usual soft terms and $D$- and $F$-term contributions.
-In this work, we take all parameters in the potential ara real.
-By using the effective potential methods epm ; epm1 ; epm2 ; epm3 ; epm4 ; epm5 ; epm6 ; epm7 ; epm8 ; epm9 ; epm10 ; epm11 ; epm12 ; epm13 ; epm14 ; epm15 ; epm16 , one can get the one-loop efffective potential:
+In this work, we take all parameters in the potential are real.
+By using the effective potential methods epm ; epm1 ; epm2 ; epm3 ; epm4 ; epm5 ; epm6 ; epm7 ; epm8 ; epm9 ; epm10 ; epm11 ; epm12 ; epm13 ; epm14 ; epm15 ; epm16 , one can get the one-loop effective potential:
 $V^{1}={{1}\over{32\pi^{2}}}\left\{\sum_{\tilde{f}}N_{f}m_{\tilde{f}}^{4}\left(\log{{m_{\tilde{f}}^{2}}\over{Q^{2}}}-{{3}\over{2}}\right)-2\sum_{f=t,b,\tau}N_{f}m_{f}^{4}\left(\log{{m_{f}^{2}}\over{Q^{2}}}-{{3}\over{2}}\right)\right\}.$ (10)
@@
 $f=t,b,\tau$ denote the third fermions and $\tilde{f}=\tilde{t}_{1,2}=\tilde{b}_{1,2}=\tilde{\tau}_{1,2}$ are the corresponding supersymmetric partners.
-Considering the one-loop effective potential, the Higgs potential can be writtrn as:
+Considering the one-loop effective potential, the Higgs potential can be written as
 $V=V^{0}+V^{1}$ (11)
-we can calculte the minimization conditions of the potential and the Higgs masses in this work.
+we can calculate the minimization conditions of the potential and the Higgs masses in this work.
 II.2 Higgs masses and CPV phases
-In the $\mu\nu$SSM, the left- & right-handed sneutrino VEVs lead to the mixing of the neutral components of the Higgs doublets with the left- &\&& right-handed sneutrinos producing an $8\times 8$ CP-even neutral scalar mass matrix, which can be seen in Refs.
+In the $\mu$-from-$\nu$SSM, the left- and right-handed sneutrino VEVs lead to the mixing of the neutral components of the Higgs doublets with the left- and right-handed sneutrinos producing an $8\times 8$ $CP$-even neutral scalar mass matrix, which can be seen in Refs.
 mnSSM1 ; mnSSM1-1 ; Zhang1 ; MASS .
-The mixing gives a rich phenomenology in the Higgs sector of the $\mu\nu$SSM mnSSM1 ; mnSSM2 ; mnSSM1-1 ; parameter-space ; mnSSM2-1 ; HZrr ; muon ; MASS ; HZr .
-Here, we note that the Higgs doublets and right-handed sneutrinos are basically decoupled from the left-handed sneutrinos MASS , so we didn’t considering the left-handed sneutrinos in the CP-even and CP-odd scalars part.
+The mixing gives a rich phenomenology in the Higgs sector of the $\mu$-from-$\nu$SSM mnSSM1 ; mnSSM2 ; mnSSM1-1 ; parameter-space ; mnSSM2-1 ; HZrr ; muon ; MASS ; HZr .
+Here, we note that the Higgs doublets and right-handed sneutrinos are basically decoupled from the left-handed sneutrinos MASS , so we did not consider the left-handed sneutrinos in the $CP$-even and $CP$-odd scalars part.
 The $CP$-even sector mix with the $CP$-odd sector, the 10 $\times$ 10 mixing matrix is defined by
 $M_{h}^{2}=\left(\begin{array}[]{cc}M_{S}^{2}&M_{SP}^{2}\\ (M_{SP}^{2})^{2}&M_{P}^{2}\\ \end{array}\right)$ (14)
-with $M_{S}^{2}$ denote the CP-even neutral scalars, $M_{P}^{2}$ is the CP-odd neutral scalars, $M_{SP}^{2}$ represents the mass submatrix for the mixing of CP-even neutral scalars and CP-odd neutral scalars.
-The mass sqared matrix $M_{h}^{2}$ can be diagonalized as:
+with $M_{S}^{2}$ denoting the $CP$-even neutral scalars, $M_{P}^{2}$ is the $CP$-odd neutral scalars, and $M_{SP}^{2}$ represents the mass submatrix for the mixing of $CP$-even neutral scalars and $CP$-odd neutral scalars.
+The mass squared matrix $M_{h}^{2}$ can be diagonalized as
 $Z_{H}M_{h}^{2}Z_{H}^{T}=m_{h}^{2}$ (15)
 with CPV in the $CP$-even and $CP$-odd scalar sector the matrix $Z_{H}$ can be complex.
-We consider the radiative correcttions in mass submatrix $M_{H}^{2}$, the radiative corrections from the third fermions ($f=t,b,\tau$) and their superpartners include the two-loop leading-log effects leading-log ; leading-log1 ; leading-log2 .
+We consider the radiative corrections in mass submatrix $M_{H}^{2}$; the radiative corrections from the third fermions ($f=t,b,\tau$) and their superpartners include the two-loop leading-log effects leading-log ; leading-log1 ; leading-log2 .
 The $CP$-even neutral scalars is given as:
 $M_{S}^{2}=\left(\begin{array}[]{cc}M_{H}^{2}&M_{Mix}^{2}\\ (M_{Mix}^{2})^{2}&M_{\tilde{\nu}_{R}^{c}}^{2}\\ \end{array}\right)$ (18)
-In detail, the mass submatrix $M_{H}^{2}$ is defind by
+In detail, the mass submatrix $M_{H}^{2}$ is defined by
 $M_{H}^{2}=\left(\begin{array}[]{cc}M_{h_{d}h_{d}}^{2}+\Delta_{11}&M_{h_{d}h_{u}}^{2}+\Delta_{12}\\ M_{h_{d}h_{u}}^{2}+\Delta_{12}&M_{h_{u}h_{u}}^{2}+\Delta_{22}\\ \end{array}\right)$ (21)
-The dominating contributions of radiative corrections $\Delta_{11},\Delta_{12}$ and $\Delta_{22}$ comes from the fermions ($t,b$) and their superpartners
+The dominating contributions of radiative corrections $\Delta_{11},~{}\Delta_{12}$ and $\Delta_{22}$ comes from the fermions ($t,b$) and their superpartners
 $\Delta_{11}=\Delta_{11}^{t}+\Delta_{11}^{b}$
@@
 $\Delta_{22}=\Delta_{22}^{t}+\Delta_{22}^{b}$ (22)
-We didn’t consider the terms containing coupling $Y_{\nu_{i}}$ and $\upsilon_{\nu_{i}}$, because these terms are very small.
-The radiative corrections from the top quark is give by HZrr ; radiative corr ; radiative corr1 ; radiative corr2 ; radiative corr3 ; radiative corr4 ; radiative corr5
-$\Delta_{11}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}{\beta}}}{{(\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu-|\mu|^{2}\cot\beta)^{2}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (23)
-$\Delta_{12}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}{\beta}}}{{|\mu|^{2}\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (24)
-×$\times\left(\ln{{m_{\tilde{t}_{1}^{2}}^{2}}\over{m_{\tilde{t}_{2}}^{2}}}+{{\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\right)$
-$\Delta_{22}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}{\beta}}}\Bigg{\{}\ln\frac{m_{\tilde{t}_{1}}^{2}m_{\tilde{t}_{2}}^{2}}{m_{t^{4}}}+{{2|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cos\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}\ln{{m_{\tilde{t}_{1}^{2}}^{2}}\over{m_{\tilde{t}_{2}}^{2}}}$ (25)
-$+{{(2|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta)^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2})^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})+{{1}\over{16\pi^{2}}}\ln{{m_{\tilde{t}_{1}}^{2}m_{\tilde{t}_{2}}^{2}}\over{m_{t}^{4}}}$
+We did not consider the terms containing coupling $Y_{\nu_{i}}$ and $\upsilon_{\nu_{i}}$, because these terms are very small.
+The radiative corrections from the top quark is given by HZrr ; radiative corr ; radiative corr1 ; radiative corr2 ; radiative corr3 ; radiative corr4 ; radiative corr5
+$\Delta_{11}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}{\beta}}}{{(\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu-|\mu|^{2}\cot\beta)^{2}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (23)
+$\Delta_{12}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}{\beta}}}{{|\mu|^{2}\cot\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (24)
+×$\times\left(\ln{{m_{\tilde{t}_{1}^{2}}^{2}}\over{m_{\tilde{t}_{2}}^{2}}}+{{\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\right)$
+$\Delta_{22}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}{\beta}}}\Bigg{\{}\ln\frac{m_{\tilde{t}_{1}}^{2}m_{\tilde{t}_{2}}^{2}}{m_{t^{4}}}+{{2|A_{t}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cos\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}\ln{{m_{\tilde{t}_{1}^{2}}^{2}}\over{m_{\tilde{t}_{2}}^{2}}}$ (25)
+$+{{(2|A_{t}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta)^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2})^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})+{{1}\over{16\pi^{2}}}\ln{{m_{\tilde{t}_{1}}^{2}m_{\tilde{t}_{2}}^{2}}\over{m_{t}^{4}}}$
 ×$\times\left({{3e^{2}m_{t}^{2}}\over{4s_{w}^{2}m_{W}^{2}}}-32\pi\alpha_{s}\right)\Bigg{[}{1\over 2}\ln{{m_{\tilde{t}_{1}}^{2}m_{\tilde{t}_{2}}^{2}}\over{m_{t}^{4}}}+{{2|A_{t}-\mu\cot\beta|^{2}}\over{m_{\tilde{t}_{1}}m_{\tilde{t}_{2}}}}\Bigg{(}1-{{(A_{t}-\mu\cot\beta)^{2}}\over{12m_{\tilde{t}_{1}}^{2}m_{\tilde{t}_{2}}^{2}}}\Bigg{)}\Bigg{]}\Bigg{\}}$
@@
 $g(m_{1}^{2},m_{2}^{2})=2-{{m_{1}^{2}+m_{2}^{2}}\over{m_{1}^{2}-m_{2}^{2}}}\ln{{m_{1}^{2}}\over{m_{2}^{2}}}$ (26)
-to save space, the mass matrix and the radiative corrections are given in the Appendix A.
-In the radiative corrections, the trilinear coupling $A_{t}=|A_{t}|e^{i\phi_{A_{t}}}$ can be complex. These seven independent phase which have been defined as:
-$\phi_{\lambda},\phi_{A_{t}},\phi_{\upsilon_{u}},\phi_{\upsilon_{d}},\phi_{\upsilon_{\nu_{i}^{c}}}$ (27)
+to save space, the mass matrix and the radiative corrections are given in the Appendix.
+In the radiative corrections, the trilinear coupling $A_{t}=|A_{t}|e^{i\phi_{A_{t}}}$ can be complex. These seven independent phase have been defined as
+$\phi_{\lambda},\phi_{A_{t}},\phi_{\upsilon_{u}},\phi_{\upsilon_{d}},\phi_{\upsilon_{\nu_{i}^{c}}}.$ (27)
 III Excess at 95 GeV
@@
 The value for $\mu_{LEP}^{bb}$ can be found in 98LEP1 ; 98LHC with the method introduced in 98LEP2 .
-$h_{1}$ is the Higgs which mass range around $\sim$ 96 GeV and $h_{2}$ is the 125 GeV Higgs boson in the following.
-In the feamework of the $\mu\nu$SSM we use $\mu_{NP}^{bb}$ to describe the signal strength, the expression for $\mu_{NP}^{bb}$ can be approximated as 98LEP3
+$h_{1}$ is the Higgs which has mass range around $\sim$ 96 GeV, and $h_{2}$ is the 125 GeV Higgs boson in the following.
+In the framework of the $\mu$-from-$\nu$ SSM we use $\mu_{NP}^{bb}$ to describe the signal strength; the expression for $\mu_{NP}^{bb}$ can be approximated as 98LEP3
 $\mu_{NP}^{bb}\!\!\!\!\!$ $={{\sigma^{\text{NP}}(Z^{*}\rightarrow Zh_{1})}\over{\sigma^{\text{SM}}(Z^{*}\rightarrow Zh_{1})}}\times{{Br^{\text{NP}}(h_{1}\rightarrow b\bar{b})}\over{Br^{\text{SM}}(h_{1}\rightarrow b\bar{b})}}$ (29)
@@
 $\approx{{|C_{h_{1}VV}|^{2}\times|C_{h_{1}d\bar{d}}|^{2}}\over{|C_{h_{1}d\bar{d}}|^{2}(Br_{h_{1}\rightarrow b\bar{b}}^{\text{SM}}+Br_{h_{1}\rightarrow\tau\bar{\tau}}^{\text{SM}})+|C_{h_{1}u\bar{u}}|^{2}(Br_{h_{1}\rightarrow g\bar{g}}^{\text{SM}}+Br_{h_{1}\rightarrow c\bar{c}}^{\text{SM}})}}$
-One can found the SM branching ratios $Br^{SM}$ in Ref.
-PDG1 , $\Gamma$ is the decay widths.
-The couplings are normalized to the SM-prediction of a Higgs boson of the same mass.
-$C_{h_{1}}$ is the coupling of $h_{1}$ and gauge boson, $C_{h_{1}u\bar{u}}$ and $C_{h_{1}d\bar{d}}$ are the couplings of $h_{1}$ and up- and down-type quarks.
-The normalized couplings are given in:
-$C_{h_{i}d\bar{d}}={{Z_{H}^{i1}}\over{\cos\beta}},\qquad C_{h_{i}u\bar{u}}={{Z_{H}^{i2}}\over{\sin\beta}},\qquad C_{h_{i}VV}=Z_{H}^{i1}\cos\beta+Z_{H}^{i2}\sin\beta,$ (30)
+One can find the SM branching ratios $Br^{SM}$ in Ref.
+PDG1 , and $\Gamma$ is the decay widths.
+The couplings are normalized to the SM prediction of a Higgs boson of the same mass.
+$C_{h_{1}}$ is the coupling of $h_{1}$ and gauge boson, and $C_{h_{1}u\bar{u}}$ and $C_{h_{1}d\bar{d}}$ are the couplings of $h_{1}$ and up- and down-type quarks.
+The normalized couplings are given as
+$C_{h_{i}d\bar{d}}={{Z_{H}^{i1}}\over{\cos\beta}},\qquad C_{h_{i}u\bar{u}}={{Z_{H}^{i2}}\over{\sin\beta}},\qquad C_{h_{i}VV}=Z_{H}^{i1}\cos\beta+Z_{H}^{i2}\sin\beta.$ (30)
 […]
-$\mu_{\gamma\gamma}^{CMS}={{\sigma(gg\rightarrow h_{1}\rightarrow\gamma\gamma)}\over{\sigma^{SM}(gg\rightarrow h_{1}\rightarrow\gamma\gamma)}}=0.6\pm 0.2,$ (31)
-Recently, ATLAS reported their new results at 95.4 GeV based on full Run 2 data set 95ATLAS , the “di-photon excess” with a signal strength of
+$\mu_{\gamma\gamma}^{CMS}={{\sigma(gg\rightarrow h_{1}\rightarrow\gamma\gamma)}\over{\sigma^{SM}(gg\rightarrow h_{1}\rightarrow\gamma\gamma)}}=0.6\pm 0.2.$ (31)
+Recently, ATLAS reported their new results at 95.4 GeV based on the full Run 2 dataset 95ATLAS , the “diphoton excess” with a signal strength of
 $\mu_{\gamma\gamma}^{ATLAS}=0.18\pm 0.1$ (32)
-Meanwhile, the corresponding CMS result for the “di-photon excess” is given by CMS4
+Meanwhile, the corresponding CMS result for the “diphoton excess” is given by CMS4
 $\mu_{\gamma\gamma}^{CMS}={{\sigma^{exp}(pp\rightarrow\phi\rightarrow\gamma\gamma)}\over{\sigma^{SM}(pp\rightarrow H_{SM}\rightarrow\gamma\gamma)}}=0.33_{-0.12}^{+0.19}.$ (33)
@@
 $\mu_{\gamma\gamma}^{exp}=\mu_{\gamma\gamma}^{ATLAS+CMS}=0.24_{-0.08}^{+0.09}.$ (34)
-In this work, the approximation of the diphoton rate of the $h_{1}$ can written as 98LEP1 ; 98LEP3 :
+In this work, the approximation of the diphoton rate of the $h_{1}$ can written as 98LEP1 ; 98LEP3
 $\mu_{NP}^{\gamma\gamma}\!\!\!\!\!$ $={{\sigma^{\text{NP}}(gg\rightarrow h_{1})}\over{\sigma^{\text{SM}}(gg\rightarrow h_{1})}}\times{{Br^{\text{NP}}(h_{1}\rightarrow\gamma\gamma)}\over{Br^{\text{SM}}(h_{1}\rightarrow\gamma\gamma)}}$ (35)
 $\approx|C_{h_{1}u\bar{u}}|^{2}\times{{\Gamma_{h_{1}\rightarrow\gamma\gamma}^{\text{NP}}}\over{\Gamma_{h_{1}\rightarrow\gamma\gamma}^{\text{SM}}}}\times{{\Gamma_{\text{tot}}^{\text{SM}}}\over{\Gamma_{\text{tot}}^{\text{NP}}}}$
-$\approx{{|C_{h_{1}u\bar{u}}|^{2}\times|C_{h_{1}\gamma\gamma}|^{2}}\over{|C_{h_{1}d\bar{d}}|^{2}(Br_{h_{1}\rightarrow b\bar{b}}^{\text{SM}}+Br_{h_{1}\rightarrow\tau\bar{\tau}}^{\text{SM}})+|C_{h_{1}u\bar{u}}|^{2}(Br_{h_{1}\rightarrow gg}^{\text{SM}}+Br_{h_{1}\rightarrow c\bar{c}}^{\text{SM}})}}$
+$\approx{{|C_{h_{1}u\bar{u}}|^{2}\times|C_{h_{1}\gamma\gamma}|^{2}}\over{|C_{h_{1}d\bar{d}}|^{2}(Br_{h_{1}\rightarrow b\bar{b}}^{\text{SM}}+Br_{h_{1}\rightarrow\tau\bar{\tau}}^{\text{SM}})+|C_{h_{1}u\bar{u}}|^{2}(Br_{h_{1}\rightarrow gg}^{\text{SM}}+Br_{h_{1}\rightarrow c\bar{c}}^{\text{SM}})}}.$
 The effective coupling $C_{h_{1}\gamma\gamma}$ can be written as 98LEP3
-$|C_{h_{1}\gamma\gamma}|^{2}={{|{4\over 3}C_{h_{1}t\bar{t}}A_{1/2}(\tau_{t})+C_{h_{1}VV}A_{1}(\tau_{W})|^{2}}\over{|{4\over 3}A_{1/2}(\tau_{t})+A_{1}(\tau_{W})|^{2}}}.$ (36)
-with $\tau_{t}={{m_{h_{1}}^{2}}\over{4m_{t}^{2}}}<1$ and $\tau_{t}={{m_{h_{1}}^{2}}\over{4m_{W}^{2}}}<1$, the from factor $A_{1/2}$ and $A_{1}$ are given by Chrr
+$|C_{h_{1}\gamma\gamma}|^{2}={{|{4\over 3}C_{h_{1}t\bar{t}}A_{1/2}(\tau_{t})+C_{h_{1}VV}A_{1}(\tau_{W})|^{2}}\over{|{4\over 3}A_{1/2}(\tau_{t})+A_{1}(\tau_{W})|^{2}}}$ (36)
+with $\tau_{t}={{m_{h_{1}}^{2}}\over{4m_{t}^{2}}}<1$ and $\tau_{t}={{m_{h_{1}}^{2}}\over{4m_{W}^{2}}}<1$. The form factors $A_{1/2}$ and $A_{1}$ are given by Chrr
 $A_{1/2}(x)\!\!\!\!\!$ $=2(x+(x-1)\arcsin^{2}\sqrt{x})x^{-2},\qquad\qquad\qquad x\leq 1$ (37)
 $A_{1}(x)\!\!\!\!\!$ $=-(2x^{2}+3x+3(2x-1)\arcsin^{2}\sqrt{x})x^{-2}\qquad x\leq 1$ (38)
-By using Eq.(29) and Eq.(35), we calculate the two signal strengths.
+By using Eqs. (29) and (35), we calculate the two signal strengths.
 IV Numerical Results
 In this section we will discuss the couplings and signal strength of 96 GeV Higgs in CPV $\mu\nu$SSM. The free parameters in our analysis will be
-$\lambda,\tan\beta,\kappa,A_{t},A_{\lambda},A_{\kappa},A_{b},\upsilon_{\nu^{c}}$ (39)
-we take $\upsilon_{\nu_{1}^{c}}=\upsilon_{\nu_{2}^{c}}=\upsilon_{\nu_{3}^{c}}$, and we have defined :
+$\lambda,\quad\tan\beta,\quad\kappa,\quad A_{t},\quad A_{\lambda},\quad A_{\kappa},\quad A_{b},\quad\upsilon_{\nu^{c}}.$ (39)
+We take $\upsilon_{\nu_{1}^{c}}=\upsilon_{\nu_{2}^{c}}=\upsilon_{\nu_{3}^{c}}$, and we have defined
 ${\kappa_{ijk}}=\kappa{\delta_{ij}}{\delta_{jk}},\quad{({A_{\kappa}}\kappa)_{ijk}}={A_{\kappa}}\kappa{\delta_{ij}}{\delta_{jk}},\quad\lambda_{i}=\lambda,$
@@
 $m_{{{\tilde{L}}_{ij}}}^{2}=m_{{\tilde{L}}}^{2}{\delta_{ij}},\quad m_{\tilde{e}_{ij}^{c}}^{2}=m_{{{\tilde{e}}^{c}}}^{2}{\delta_{ij}},\quad\upsilon_{\nu_{i}^{c}}=\upsilon_{\nu^{c}},$ (40)
-where $i,j=1,2,3$
+where $i,j=1,2,3$.
 IV.1 Mass and coupling
-Figure 1: Values for $m_{h_{2}}$ versus $m_{h_{2}}$ in (a), the normalized couplings $|C_{h_{2}VV}|$ versus $|C_{h_{1}VV}|$ in (b). The input parameters in Tab. 1.
-[…]
-Figure 2: Values for $m_{h_{1}}$ versus the parameters $\upsilon_{\nu^{c}}$ in (a) and $\kappa$ in (c), values for $m_{h_{2}}$ versus the parameters $\upsilon_{\nu^{c}}$ in (b) and $\kappa$ in (d).
-We take $\tan\beta=5$ and other parameters in Tab.
-1.
-In Fig.2, the left column, we can see that $m_{h_{1}}$ is very sensitive to $\upsilon_{\nu^{c}}$.
-As $\upsilon_{\nu^{c}}$ becomes larger, $m_{h_{1}}$ will rapidly become smaller, on the contrary, $m_{h_{2}}$ will slowly become larger.
+Figure 1: Values for $m_{h_{2}}$ versus $m_{h_{2}}$ in (a) and the normalized couplings $|C_{h_{2}VV}|$ versus $|C_{h_{1}VV}|$ in (b). The input parameters are in Table 1.
+In Fig.
+1 (a), we can see that, when $h_{1}$ is near 96 GeV, $h_{2}$ is closer to 125 GeV with the increase of $\tan\beta$.
+Although both $h_{1}$ and $h_{2}$ can conform to the experimental mass range in our parameter space, if we assume a theory uncertainty of up to 3 GeV 98LEP3 , the parameter range will be larger.
+One can clear see that $|C_{h_{2}VV}|$ is much larger than $|C_{h_{1}VV}|$ in Fig.
+1 (b); the LHC measurements of the SM-like Higgs bosons couplings to fermions and massive gauge bosons are still not very precise98LEP3 ; LHC-couplings .
+If in the future some collider can measure these couplings to the percent level, then we can choose a more reasonable parameter space.
+Figure 2: Values for $m_{h_{1}}$ versus the parameters $\upsilon_{\nu^{c}}$ in (a) and $\kappa$ in (c) and values for $m_{h_{2}}$ versus the parameters $\upsilon_{\nu^{c}}$ in (b) and $\kappa$ in (d).
+We take $\tan\beta=5$ and other parameters in Table 1.
+In Fig.2, left column, we can see that $m_{h_{1}}$ is very sensitive to $\upsilon_{\nu^{c}}$.
+As $\upsilon_{\nu^{c}}$ becomes larger, $m_{h_{1}}$ will rapidly become smaller; on the contrary, $m_{h_{2}}$ will slowly become larger.
 In order for $m_{h_{1}}$ to be around 95.4 GeV, the value of $\upsilon_{\nu^{c}}$ will not have a wide range.
 Similar to $\upsilon_{\nu^{c}}$, $m_{h_{1}}$ is also very sensitive to $\kappa$, and with the increase of $\lambda$, a larger $\kappa$ value can be taken to keep $m_{h_{1}}$ near 95.4 GeV.
-On the contrary, $\kappa$ has little effect on $m_{h_{2}}$, especially when $\lambda$=0.085 or 0.087, $m_{h_{2}}$ will be very stable.
+On the contrary, $\kappa$ has little effect on $m_{h_{2}}$; especially when $\lambda$=0.085 or 0.087, $m_{h_{2}}$ will be very stable.
 $\qquad A_{\lambda}\qquad$ $\qquad\lambda\qquad$ $\qquad A_{t}\qquad$ $\qquad\upsilon_{\nu^{c}}\qquad$ $\qquad\kappa\qquad$ $\qquad A_{\kappa}\qquad$
-920 [0.08;0.09] 1660 1000 0.042 -370
+920 [0.08;0.09] 1660 1000 0.042 --370
 $\qquad A_{b}\qquad$ $\qquad\phi_{\lambda}/\pi\qquad$ $\qquad\phi_{A_{t}}/\pi\qquad$ $\qquad\phi_{v_{u}}/\pi\qquad$ $\qquad\phi_{v_{d}}/\pi\qquad$ $\qquad\phi_{\upsilon_{\nu^{c}}}/\pi\qquad$
-700 -0.020 -0.200 0.380 -0.095 -1
+700 $-0.020$-0.200 0.380 --0.095 --1
 Table 1: Input parameters to fit the LEP and the CMS excesses. All dimensionful parameters are given in GeV.
-Figure 3: The component of $h_{1}$ (a) and $h_{2}$ (b). The input parameters in Tab.1, and we take $\tan\beta=4.31$.
+Figure 3: The component of $h_{1}$ (a) and $h_{2}$ (b). The input parameters are in Table 1, and we take $\tan\beta=4.31$.
 In Fig.
-3, we have showed the component of $h_{1}$ and $h_{2}$, for $h_{1}$, CP-odd $H_{u}$ component and CP-even right-handed sneutrinos component are the main components, and with the increase of $\lambda$, CP-odd $H_{d}$ component will become smaller, while the CP-even right-handed sneutrino component will become larger.
+3, we have showed the component of $h_{1}$ and $h_{2}$; for $h_{1}$, $CP$-odd $H_{u}$ component and $CP$-even right-handed sneutrinos component are the main components, and with the increase of $\lambda$, $CP$-odd $H_{d}$ component will become smaller, while the $CP$-even right-handed sneutrino component will become larger.
 While for $h_{2}$, $CP$-even $H_{u}$ is the main component, as $\lambda$ increases, the $CP$-even $H_{u}$ component will increase, the $CP$-even $H_{d}$ component will gradually decrease, and the $CP$-odd $H_{d}$ component will gradually increase.
-Figure 4: The left column is the correlation of $h_{1}$ and the CPV phases, the grey area represents the experimental error of 3 GeV.
-The right column is the correlation of $h_{2}$ and the CPV phases, the grey area is the theory uncertainty of 1 GeV.
+Figure 4: The correlation of $h_{1}$ and the CPV phases in (a) (c) (e) (g); the gray area represents the experimental error of 3 GeV.
+The correlation of $h_{2}$ and the CPV phases in (b) (d) (f) (h); the gray area is the theory uncertainty of 1 GeV.
 Then, in Fig.
 4, we analyze the correlation between the Higgs masses and the CPV phases.
-For the first row, we take $\lambda=0.09,\tan\beta=6,\kappa=0.315,$ for the second row, $\lambda=0.095,\tan\beta=4.5,\kappa=0.315,$ for the third row, $\lambda=0.098,\tan\beta=4.3,\kappa=0.32$, for the last row, $\lambda=0.085,\tan\beta=3,\kappa=0.04$.
-Other parameter keep same as Tab.1.
-We should remark that there are many CPV phase values that can be constrained by the Higgs masses of Fig.7 (a), and here we only select one of them.
-The signal strengths and the Higgs masses do not increase or decrease all the time, but oscillates like a sine function.
+For the first row, we take $\lambda=0.09,\tan\beta=6,$ and $\kappa=0.315,$ for the second row, $\lambda=0.095,\tan\beta=4.5,$ and $\kappa=0.315,$ for the third row, $\lambda=0.098,\tan\beta=4.3,$ and $\kappa=0.32$, for the last row, $\lambda=0.085,\tan\beta=3,$ and $\kappa=0.04$.
+Other parameters remain the same as Table 1.
+We should remark that there are many CPV phase values that can be constrained by the Higgs masses in Fig.7 (a), and here we select only one of them.
+The signal strengths and the Higgs masses do not increase or decrease all the time, but oscillate like a sine function.
 In Fig.
-4, the first row, when $\phi_{\lambda}$ grown from 4 to 5, the mass range of the “di-photon excess” can fit the experimental constraints, the SM-like Higgs mass can be kept around 125 GeV.
+4, the first row, when $\phi_{\lambda}$ grows from 4 to 5, the mass range of the “diphoton excess” can fit the experimental constraints, and the SM-like Higgs mass can be kept around 125 GeV.
 […]
 IV.2 Signal strengths
-Figure 5: The left column shows that the signal strength $\mu_{NP}^{bb}$ varies with $\upsilon_{\nu^{c}}$ (a) and $\kappa$ (c), the right column shows that the signal strength $\mu_{NP}^{\gamma\gamma}$ varies with $\upsilon_{\nu^{c}}$ (b) and $\kappa$ (d).
-We take $\lambda=0.09$ and other parameters in Tab.
-1.
+Figure 5: The left column shows that the signal strength $\mu_{NP}^{bb}$ varies with $\upsilon_{\nu^{c}}$ (a) and $\kappa$ (c); the right column shows that the signal strength $\mu_{NP}^{\gamma\gamma}$ varies with $\upsilon_{\nu^{c}}$ (b) and $\kappa$ (d).
+We take $\lambda=0.09$ and other parameters as in Table 1.
 […]
-The effects of CPV phases is also very obvious.
+The effect of CPV phases is also very obvious.
 Here, we take $\phi_{A_{t}}$ as an example, we take $\phi_{A_{t}}$ from $-\pi$ to $\pi$.
-In the first row of Fig.
+In the first row in Fig.
 6, we can see that the minimum value of $h_{1}$ appears to be around 96.2 GeV and varies periodically with $\phi_{A_{t}}$.
@@
 Meanwhile, the highest point of $h_{2}$ is at 125.3 GeV, with a maximum value 1.5GeV larger than the minimum value.
-The impact of $\phi_{A_{t}}$ on the normalized couplings are relatively small.
-In the second row of Fig.
+The impact of $\phi_{A_{t}}$ on the normalized couplings is relatively small.
+In the second row in Fig.
 6, we can see that the peak of $|C_{h_{1}VV}|$ is close to 0.313, and the highest point of $|C_{h_{2}VV}|$ is slightly less than 0.856.
-This implies that if the ”95.4 excess” is a new particle, then its coupling with gauge bosons should be much smaller than the coupling of the SM-like Higgs with gauge bosons.
-Figure 6: Higgs masses, couplings, and signal strengths are sensitive to CPV phases. Here, we take $\tan\beta$=4.5, $\lambda$=0.086, $\phi_{A_{t}}=[-\pi,\pi]$, other parameters are the same as Tab. 1.
+This implies that, if the ”95.4 excess” is a new particle, then its coupling with gauge bosons should be much smaller than the coupling of the SM-like Higgs with gauge bosons.
+Figure 6: The Higgs masses vary with the phase $\phi_{A_{t}}$ in (a)-(b); the correlation of couplings and the phase $\phi_{A_{t}}$ in (c)-(d); and (e)-(f) are the correlation of signal strengths and the phase $\phi_{A_{t}}$.
+Here, we take $\tan\beta$=4.5, $\lambda$=0.086, and $\phi_{A_{t}}=[-\pi,\pi]$, and other parameters are the same as in Table 1.
 […]
 $\qquad\lambda\qquad$ $\qquad\tan\beta\qquad$ $\qquad\kappa\qquad$ $\qquad\upsilon_{\nu^{c}}\qquad$ $\qquad A_{\lambda}\qquad$ $\qquad A_{t}\qquad$ $\qquad A_{\kappa}\qquad$
-$[0.087;0.090]$ [2.4;4.5] [0.040;0.045] [950;1050] [915;920] [1600;1700] [-375;-370]
+$[0.087;0.090]$ [2.4;4.5] [0.040;0.045] [950;1050] [915;920] [1600;1700] [--375;--370]
 $\qquad A_{b}\qquad$ $\qquad\phi_{\lambda}/\pi\qquad$ $\qquad\phi_{A_{t}}/\pi\qquad$ $\qquad\phi_{v_{u}}/\pi\qquad$ $\qquad\phi_{v_{d}}/\pi\qquad$ $\qquad\phi_{\upsilon_{\nu^{c}}}/\pi\qquad$ $\qquad M_{2}\qquad$
-$[695;710]$ [-0.024;-0.022] [-0.4;-0.1] [0.382;0.385] [-0.095;-0.090] [-1;-0.9] 800
+$[695;710]$ [--0.024;--0.022] [--0.4;--0.1] [0.382;0.385] [--0.095;--0.090] [--1;--0.9] 800
 Table 2: The parameter space of the random scan plot. All dimensionful parameters are given in GeV.
 […]
-Let us remark that in the random scan plots, we must first ensure that the mass of $h_{2}$ can conform to the experimental constraints, because the mass determination for SM-like Higgs is already very accurate.
+Let us remark that, in the random scan plots, we must first ensure that the mass of $h_{2}$ can conform to the experimental constraints, because the mass determination for SM-like Higgs is already very accurate.
 In Fig.7 (a), we chose 1 GeV theory uncertainty for the SM-like Higgs.
-For $h_{1}$, we can choose the theory uncertainty up to 3 GeV, because CMS only shows that there is a 2.8 $\sigma$ excess at 95.3 GeV CMS1 ; CMS5 ; CMS9 , which is not exactly observed.
-In Fig 7 (b), the $(\mu_{NP}^{\gamma\gamma},\mu_{NP}^{bb})$ plane, we can clearly see that with the parameter space of Tab.2, most points can explain the di-photon excess and the $b\bar{b}$ excess.
+For $h_{1}$, we can choose the theory uncertainty up to 3 GeV, because CMS shows only that there is a 2.8 $\sigma$ excess at 95.3 GeV CMS1 ; CMS5 ; CMS9 , which is not exactly observed.
+In Fig 7.
+(b), the $(\mu_{NP}^{\gamma\gamma},\mu_{NP}^{bb})$ plane, we can clearly see that with the parameter space of Table 2, most points can explain the dphoton excess and the $b\bar{b}$ excess.
 V Conclusion
-In this paper we introduced CPV in the $\mu\nu$SSM, which lead CP-even Higgs sector mixed with CP-odd Higgs sector.
-We also analyzed a excess in the di-photon decay mode at $\sim$ 95 GeV as reported by ATLAS and CMS, together with a $\sim$ 2 $\sigma$ excess at LEP in the same mass range.
+In this paper we introduced CPV in the $\mu$-from-$\nu$SSM, which leads to $CP$-even Higgs sector mixed with $CP$-odd Higgs sector.
+We also analyzed an excess in the diphoton decay mode at $\sim$ 95 GeV as reported by ATLAS and CMS, together with a $\sim$ 2 $\sigma$ excess at LEP in the same mass range.
 The mixing and CPV are used to produce the lightest Higgs boson mass around 95 GeV, and the next-to-lightest Higgs boson mass around 125 GeV is the so-called SM-like Higgs.
 The lightest Higgs boson can explain an excess of $\gamma\gamma$ events at $\sim$ 96 GeV as reported by CMS.
-In the numerical part, we find a suitable parameter spaces, based on which we show the properties of the lightest and next-to-lightest Higgs boson.
+In the numerical part, we find a suitable parameter space, based on which we show the properties of the lightest and next-to-lightest Higgs boson.
 We found it very easy get both $\mu_{NP}^{bb}$ and $\mu_{NP}^{\gamma\gamma}$ to reach the central values of their respective experiments at the same time.
-We also analyze the influence of relevant parameters and CPV phases on the signal strengths or Higgs masses, and improve the signal strengths as much as possible while ensuring that the SM-like Higgs meets the experimental constraints and the “di-photon excess” is around 95 GeV.
+We also analyze the influence of relevant parameters and CPV phases on the signal strengths or Higgs masses, and improve the signal strengths as much as possible while ensuring that the SM-like Higgs meets the experimental constraints and the “diphoton excess” is around 95 GeV.
 Acknowledgements.
@@
 12235008, No.
-11705045, No.
-11535002, Hebei Natural Science Foundation with Grant No.
-A2022201017, No.
+11705045, and No.
+11535002, Hebei Natural Science Foundation with Grants No.
+A2022201017 and No.
 A2023201041, Natural Science Foundation of Guangxi Autonomous Region with Grant No.
-2022GXNSFDA035068, the youth top-notch talent support program of the Hebei Province.
-Appendix A Msss matyices
-The CP-even neutral scalars have the composition $S^{T}=(h_{d},h_{u},(\tilde{\nu}_{i}^{c})^{\Re})$, one can write the mass matrix $M_{S}^{2}$ as:
-$M_{S}^{2}=\left(\begin{array}[]{ccccc}M_{h_{d}h_{d}}^{2}&M_{h_{d}h_{u}}^{2}&M_{h_{d}(\tilde{\nu}_{1}^{c})^{\Re}}^{2}&M_{h_{d}(\tilde{\nu}_{2}^{c})^{\Re}}^{2}&M_{h_{d}(\tilde{\nu}_{3}^{c})^{\Re}}^{2}\\ M_{h_{u}h_{d}}^{2}&M_{h_{u}h_{u}}^{2}&M_{h_{u}(\tilde{\nu}_{1}^{c})^{\Re}}^{2}&M_{h_{u}(\tilde{\nu}_{2}^{c})^{\Re}}^{2}&M_{h_{u}(\tilde{\nu}_{3}^{c})^{\Re}}^{2}\\ M_{(\tilde{\nu}_{1}^{c})^{\Re}h_{d}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}h_{u}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Re}}^{2}\\ M_{(\tilde{\nu}_{2}^{c})^{\Re}h_{d}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}h_{u}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Re}}^{2}\\ M_{(\tilde{\nu}_{3}^{c})^{\Re}h_{d}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}h_{u}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Re}}^{2}\\ \end{array}\right)$ (46)
-$M_{h_{d}h_{d}}^{2}\!\!\!\!\!$ $={{G^{2}}\over{2}}v_{d}^{2}+\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\tan\beta$ (47)
-$+\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\tan\beta+\Delta_{11},$
-$M_{h_{d}h_{u}}^{2}\!\!\!\!\!$ $=-{{G^{2}}\over{2}}v_{d}v_{u}-\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}+2|\lambda_{i}|^{2}v_{d}v_{u}$ (48)
-$-\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}+\Delta_{12},$
-$M_{h_{u}h_{u}}^{2}\!\!\!\!\!$ $={{G^{2}}\over{2}}v_{u}^{2}+\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\cot\beta$ (49)
-$+\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\cot\beta+\Delta_{22},$
-$M_{h_{d}(\tilde{\nu}_{i}^{c})^{\Re}}^{2}\!\!\!\!\!$ $=-\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}+2\lambda_{i}\lambda_{j}^{*}v_{d}\upsilon_{\nu_{j}^{c}}$ (50)
-$-2\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{u}\upsilon_{\nu_{j}^{c}}+\Delta_{1(2+i)},$
-$M_{h_{u}(\tilde{\nu}_{i}^{c})^{\Re}}^{2}\!\!\!\!\!$ $=-\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}+2\lambda_{i}\lambda_{j}^{*}v_{u}\upsilon_{\nu_{j}^{c}}$ (51)
-$-2\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}\upsilon_{\nu_{j}^{c}}+\Delta_{2(2+i)},$
-$M_{(\tilde{\nu}_{i}^{c})^{\Re}(\tilde{\nu}_{i}^{c})^{\Re}}^{2}\!\!\!\!\!$ $=(A_{\kappa}\kappa)_{ijk}\Re(e^{i\phi_{\upsilon_{\nu_{k}^{c}}}}e^{i\phi_{\upsilon_{\nu_{k}^{c}}}}e^{i\phi_{\upsilon_{\nu_{k}^{c}}}})\upsilon_{\nu_{k}^{c}}+\Delta_{(2+i)(2+j)}$ (52)
-$+\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}{{v_{d}v_{u}}\over{\upsilon_{\nu_{j}^{c}}}},$
-$M_{(\tilde{\nu}_{i}^{c})^{\Re}(\tilde{\nu}_{j}^{c})^{\Re}}^{2}\!\!\!\!\!$ $=|\lambda_{i}|^{2}(v_{d}^{2}+v_{u}^{2})+\Delta_{(2+i)(2+j)}.$ (53)
-The rediative corrections from the top quark and bottom quark and their corresponding supersymmetric partners, the corrections in the mass matrix can be expressed as:
+2022GXNSFDA035068, and the youth top-notch talent support program of the Hebei Province.
+Appendix A mass matrices
+The $CP$-even neutral scalars have the composition $S^{T}=(h_{d},h_{u},(\tilde{\nu}_{i}^{c})^{\text{Re}})$, and one can write the mass matrix $M_{S}^{2}$ as
+$M_{S}^{2}=\left(\begin{array}[]{ccccc}M_{h_{d}h_{d}}^{2}&M_{h_{d}h_{u}}^{2}&M_{h_{d}(\tilde{\nu}_{1}^{c})^{\text{Re}}}^{2}&M_{h_{d}(\tilde{\nu}_{2}^{c})^{\text{Re}}}^{2}&M_{h_{d}(\tilde{\nu}_{3}^{c})^{\text{Re}}}^{2}\\ M_{h_{u}h_{d}}^{2}&M_{h_{u}h_{u}}^{2}&M_{h_{u}(\tilde{\nu}_{1}^{c})^{\text{Re}}}^{2}&M_{h_{u}(\tilde{\nu}_{2}^{c})^{\text{Re}}}^{2}&M_{h_{u}(\tilde{\nu}_{3}^{c})^{\text{Re}}}^{2}\\ M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}h_{d}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}h_{u}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}(\tilde{\nu}_{1}^{c})^{\text{Re}}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}(\tilde{\nu}_{2}^{c})^{\text{Re}}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}(\tilde{\nu}_{3}^{c})^{\text{Re}}}^{2}\\ M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}h_{d}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}h_{u}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}(\tilde{\nu}_{1}^{c})^{\text{Re}}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}(\tilde{\nu}_{2}^{c})^{\text{Re}}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}(\tilde{\nu}_{3}^{c})^{\text{Re}}}^{2}\\ M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}h_{d}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}h_{u}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}(\tilde{\nu}_{1}^{c})^{\text{Re}}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}(\tilde{\nu}_{2}^{c})^{\text{Re}}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}(\tilde{\nu}_{3}^{c})^{\text{Re}}}^{2}\\ \end{array}\right)$ (46)
+$M_{h_{d}h_{d}}^{2}\!\!\!\!\!$ $={{G^{2}}\over{2}}v_{d}^{2}+\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\tan\beta$ (47)
+$+\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\tan\beta+\Delta_{11},$
+$M_{h_{d}h_{u}}^{2}\!\!\!\!\!$ $=-{{G^{2}}\over{2}}v_{d}v_{u}-\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}+2|\lambda_{i}|^{2}v_{d}v_{u}$ (48)
+$-\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}+\Delta_{12},$
+$M_{h_{u}h_{u}}^{2}\!\!\!\!\!$ $={{G^{2}}\over{2}}v_{u}^{2}+\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\cot\beta$ (49)
+$+\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\cot\beta+\Delta_{22},$
+$M_{h_{d}(\tilde{\nu}_{i}^{c})^{\text{Re}}}^{2}\!\!\!\!\!$ $=-\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}+2\lambda_{i}\lambda_{j}^{*}v_{d}\upsilon_{\nu_{j}^{c}}$ (50)
+$-2\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{u}\upsilon_{\nu_{j}^{c}}+\Delta_{1(2+i)},$
+$M_{h_{u}(\tilde{\nu}_{i}^{c})^{\text{Re}}}^{2}\!\!\!\!\!$ $=-\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}+2\lambda_{i}\lambda_{j}^{*}v_{u}\upsilon_{\nu_{j}^{c}}$ (51)
+$-2\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}\upsilon_{\nu_{j}^{c}}+\Delta_{2(2+i)},$
+$M_{(\tilde{\nu}_{i}^{c})^{\text{Re}}(\tilde{\nu}_{i}^{c})^{\text{Re}}}^{2}\!\!\!\!\!$ $=(A_{\kappa}\kappa)_{ijk}\text{Re}(e^{i\phi_{\upsilon_{\nu_{k}^{c}}}}e^{i\phi_{\upsilon_{\nu_{k}^{c}}}}e^{i\phi_{\upsilon_{\nu_{k}^{c}}}})\upsilon_{\nu_{k}^{c}}+\Delta_{(2+i)(2+j)}$ (52)
+$+\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}{{v_{d}v_{u}}\over{\upsilon_{\nu_{j}^{c}}}},$
+$M_{(\tilde{\nu}_{i}^{c})^{\text{Re}}(\tilde{\nu}_{j}^{c})^{\text{Re}}}^{2}\!\!\!\!\!$ $=|\lambda_{i}|^{2}(v_{d}^{2}+v_{u}^{2})+\Delta_{(2+i)(2+j)}.$ (53)
+The radiative corrections from the top quark and bottom quark and their corresponding supersymmetric partners, the corrections in the mass matrix, can be expressed as
 $\Delta_{ab}=\Delta_{ab}^{t}+\Delta_{ab}^{b}$ (54)
-$\Delta_{1(2+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}\Bigg{\{}{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (55)
-×$\times{{|\mu|^{2}\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
-$\Delta_{2(2+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}\Bigg{\{}{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}\Bigg{\}}$ (56)
-×$\times\Bigg{\{}\ln{{m_{\tilde{t}_{1}}^{2}}\over{m_{\tilde{t}_{2}}^{2}}}+{{|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
-$\Delta_{(2+i)(2+j)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}\Bigg{\{}{{({1\over 2}\upsilon_{\nu_{j}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d})^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2})^{2}}}$ (57)
+$\Delta_{1(2+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}\Bigg{\{}{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\cot\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (55)
+×$\times{{|\mu|^{2}\cot\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
+$\Delta_{2(2+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}\Bigg{\{}{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\cot\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}\Bigg{\}}$ (56)
+×$\times\Bigg{\{}\ln{{m_{\tilde{t}_{1}}^{2}}\over{m_{\tilde{t}_{2}}^{2}}}+{{|A_{t}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
+$\Delta_{(2+i)(2+j)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}\Bigg{\{}{{({1\over 2}\upsilon_{\nu_{j}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\cot\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d})^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2})^{2}}}$ (57)
 ×$\times g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}}$
-$\Delta_{11}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}\ln{{m_{\tilde{b}_{1}}^{2}m_{\tilde{b}_{2}}^{2}}\over{m_{b}^{4}}}+{{2|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\ln{{m_{\tilde{b}_{1}}^{2}}\over{m_{\tilde{b}_{2}}^{2}}}$ (58)
-$+{{(|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
-$\Delta_{12}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$ (59)
-×$\times\Bigg{\{}\ln{{m_{\tilde{b}_{1}}^{2}}\over{m_{\tilde{b}_{2}}^{2}}}+{{|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
-$\Delta_{22}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{(|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),\Bigg{\}}$ (60)
-$\Delta_{1(2+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$ (61)
-×$\times\Bigg{\{}\ln{{m_{\tilde{b}_{1}}^{2}}\over{m_{\tilde{b}_{2}}^{2}}}+{{|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
-$\Delta_{2(2+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (62)
-×$\times{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$
-$\Delta_{(2+i)(2+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{({1\over 2}\upsilon_{\nu_{j}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u})^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}$ (63)
+$\Delta_{11}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}\ln{{m_{\tilde{b}_{1}}^{2}m_{\tilde{b}_{2}}^{2}}\over{m_{b}^{4}}}+{{2|A_{b}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\ln{{m_{\tilde{b}_{1}}^{2}}\over{m_{\tilde{b}_{2}}^{2}}}$ (58)
+$+{{(|A_{b}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
+$\Delta_{12}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{|\mu|^{2}\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$ (59)
+×$\times\Bigg{\{}\ln{{m_{\tilde{b}_{1}}^{2}}\over{m_{\tilde{b}_{2}}^{2}}}+{{|A_{b}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
+$\Delta_{22}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{(|\mu|^{2}\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),\Bigg{\}}$ (60)
+$\Delta_{1(2+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$ (61)
+×$\times\Bigg{\{}\ln{{m_{\tilde{b}_{1}}^{2}}\over{m_{\tilde{b}_{2}}^{2}}}+{{|A_{b}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
+$\Delta_{2(2+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{|\mu|^{2}\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (62)
+×$\times{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$
+$\Delta_{(2+i)(2+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{({1\over 2}\upsilon_{\nu_{j}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u})^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}$ (63)
 ×$\times g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}).$
-In the same way, the CP-odd neutral scalars mass matrix is that:
-$M_{P}^{2}=\left(\begin{array}[]{ccccc}M_{\sigma_{d}\sigma_{d}}^{2}&M_{\sigma_{d}\sigma_{u}}^{2}&M_{\sigma_{d}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{\sigma_{d}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{\sigma_{d}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{\sigma_{u}\sigma_{d}}^{2}&M_{\sigma_{u}\sigma_{u}}^{2}&M_{\sigma_{u}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{\sigma_{u}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{\sigma_{u}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{1}^{c})^{\Im}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Im}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Im}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Im}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Im}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{2}^{c})^{\Im}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Im}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Im}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Im}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Im}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{3}^{c})^{\Im}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Im}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Im}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Im}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Im}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ \end{array}\right)$ (69)
-$M_{\sigma_{d}\sigma_{d}}^{2}\!\!\!\!\!$ $=\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\tan\beta$ (70)
-$+\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\tan\beta+\Delta_{66},$
-$M_{\sigma_{d}\sigma_{u}}^{2}\!\!\!\!\!$ $=-\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}$ (71)
-$+\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}+\Delta_{67},$
-$M_{\sigma_{u}\sigma_{u}}^{2}\!\!\!\!\!$ $=\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\cot\beta$ (72)
-$+\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\cot\beta+\Delta_{77},$
-$M_{\sigma_{d}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}$ (73)
-$+2\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{u}\upsilon_{\nu_{i}^{c}}+\Delta_{6(7+i)},$
-$M_{\sigma_{u}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=-\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}$ (74)
-$-2\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}\upsilon_{\nu_{i}^{c}}+\Delta_{7(7+i)},$
-$M_{(\tilde{\nu}_{i}^{c})^{\Im}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}{{v_{d}v_{u}}\over{\upsilon_{\nu_{i}^{c}}}}-(A_{\kappa}\kappa)_{ijk}\Re(e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})\upsilon_{\nu_{k}^{c}}$ (75)
-$+4\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}v_{u}+\Delta_{(7+i)(7+i)},$
-$M_{(\tilde{\nu}_{i}^{c})^{\Im}(\tilde{\nu}_{j}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=|\lambda_{i}|^{2}(v_{u}^{2}+v_{d}^{2})+\Delta_{(7+i)(7+i)}.$ (76)
-$\Delta_{66}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{(-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu)^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}})^{2}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$ (77)
-$\Delta_{67}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (78)
-×$\times{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
-$\Delta_{77}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{(\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta)^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2})^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$ (79)
-$\Delta_{6(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (80)
-×$\times{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
-$\Delta_{7(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (81)
-×$\times{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
-$\Delta_{(7+i)(7+j)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{(-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d})^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}})^{2}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (82)
-$\Delta_{66}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(-\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}})^{2}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (83)
-$\Delta_{67}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-(\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu)^{2}\tan\beta}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (84)
-$\Delta_{77}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (85)
-$\Delta_{6(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-(\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b})^{2}\mu\tan\beta\lambda_{i}v_{u}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (86)
-$\Delta_{7(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b})^{2}\mu\lambda_{i}v_{u}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (87)
-$\Delta_{(7+i)(7+j)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u})^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}).$ (88)
+In the same way, the $CP$-odd neutral scalars mass matrix is that
+$M_{P}^{2}=\left(\begin{array}[]{ccccc}M_{\sigma_{d}\sigma_{d}}^{2}&M_{\sigma_{d}\sigma_{u}}^{2}&M_{\sigma_{d}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{\sigma_{d}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{\sigma_{d}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ M_{\sigma_{u}\sigma_{d}}^{2}&M_{\sigma_{u}\sigma_{u}}^{2}&M_{\sigma_{u}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{\sigma_{u}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{\sigma_{u}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ M_{(\tilde{\nu}_{1}^{c})^{\text{Im}}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Im}}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Im}}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Im}}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Im}}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ M_{(\tilde{\nu}_{2}^{c})^{\text{Im}}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Im}}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Im}}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Im}}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Im}}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ M_{(\tilde{\nu}_{3}^{c})^{\text{Im}}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Im}}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Im}}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Im}}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Im}}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ \end{array}\right)$ (69)
+$M_{\sigma_{d}\sigma_{d}}^{2}\!\!\!\!\!$ $=\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\tan\beta$ (70)
+$+\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\tan\beta+\Delta_{66},$
+$M_{\sigma_{d}\sigma_{u}}^{2}\!\!\!\!\!$ $=-\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}$ (71)
+$+\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}+\Delta_{67},$
+$M_{\sigma_{u}\sigma_{u}}^{2}\!\!\!\!\!$ $=\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\cot\beta$ (72)
+$+\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\cot\beta+\Delta_{77},$
+$M_{\sigma_{d}(\tilde{\nu}_{i}^{c})^{\text{Im}}}^{2}\!\!\!\!\!$ $=\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}$ (73)
+$+2\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{u}\upsilon_{\nu_{i}^{c}}+\Delta_{6(7+i)},$
+$M_{\sigma_{u}(\tilde{\nu}_{i}^{c})^{\text{Im}}}^{2}\!\!\!\!\!$ $=-\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}$ (74)
+$-2\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}\upsilon_{\nu_{i}^{c}}+\Delta_{7(7+i)},$
+$M_{(\tilde{\nu}_{i}^{c})^{\text{Im}}(\tilde{\nu}_{i}^{c})^{\text{Im}}}^{2}\!\!\!\!\!$ $=\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}{{v_{d}v_{u}}\over{\upsilon_{\nu_{i}^{c}}}}-(A_{\kappa}\kappa)_{ijk}\text{Re}(e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})\upsilon_{\nu_{k}^{c}}$ (75)
+$+4\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}v_{u}+\Delta_{(7+i)(7+i)},$
+$M_{(\tilde{\nu}_{i}^{c})^{\text{Im}}(\tilde{\nu}_{j}^{c})^{\text{Im}}}^{2}\!\!\!\!\!$ $=|\lambda_{i}|^{2}(v_{u}^{2}+v_{d}^{2})+\Delta_{(7+i)(7+i)},$ (76)
+$\Delta_{66}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{(-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu)^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}})^{2}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$ (77)
+$\Delta_{67}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (78)
+×$\times{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
+$\Delta_{77}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{(\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta)^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2})^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$ (79)
+$\Delta_{6(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (80)
+×$\times{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
+$\Delta_{7(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (81)
+×$\times{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
+$\Delta_{(7+i)(7+j)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{(-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d})^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}})^{2}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (82)
+$\Delta_{66}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}})^{2}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (83)
+$\Delta_{67}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-(\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu)^{2}\tan\beta}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (84)
+$\Delta_{77}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (85)
+$\Delta_{6(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-(\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b})^{2}\mu\tan\beta\lambda_{i}v_{u}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (86)
+$\Delta_{7(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b})^{2}\mu\lambda_{i}v_{u}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (87)
+$\Delta_{(7+i)(7+j)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u})^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}).$ (88)
 The mix mass submatrix $M_{SP}^{2}$
-$M_{SP}^{2}=\left(\begin{array}[]{ccccc}M_{h_{d}\sigma_{d}}^{2}&M_{h_{d}\sigma_{u}}^{2}&M_{h_{d}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{h_{d}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{h_{d}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{h_{u}\sigma_{d}}^{2}&M_{h_{u}\sigma_{u}}^{2}&M_{h_{u}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{h_{u}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{h_{u}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{1}^{c})^{\Re}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{2}^{c})^{\Re}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{3}^{c})^{\Re}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ \end{array}\right)$ (94)
+$M_{SP}^{2}=\left(\begin{array}[]{ccccc}M_{h_{d}\sigma_{d}}^{2}&M_{h_{d}\sigma_{u}}^{2}&M_{h_{d}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{h_{d}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{h_{d}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ M_{h_{u}\sigma_{d}}^{2}&M_{h_{u}\sigma_{u}}^{2}&M_{h_{u}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{h_{u}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{h_{u}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ \end{array}\right)$ (94)
 $M_{h_{d}\sigma_{d}}^{2}\!\!\!\!\!$ $=\Delta_{16},$ (95)
@@
 $M_{h_{u}\sigma_{u}}^{2}\!\!\!\!\!$ $=\Delta_{27},$ (98)
-$M_{h_{d}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=-\Im(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}+\Delta_{1(7+i)},$ (99)
-$M_{h_{u}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=-\Im(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}+\Delta_{2(7+i)},$ (100)
-$M_{(\tilde{\nu}_{i}^{c})^{\Re}\sigma_{d}}^{2}\!\!\!\!\!$ $=-\Im(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}+\Delta_{(2+i)6},$ (101)
-$M_{(\tilde{\nu}_{i}^{c})^{\Re}\sigma_{u}}^{2}\!\!\!\!\!$ $=\Im(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}+\Delta_{(2+i)7},$ (102)
-$M_{(\tilde{\nu}_{i}^{c})^{\Re}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=2\Im(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}v_{u}$ (103)
-$-(A_{\kappa}\kappa)_{i}\Re(e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})\upsilon_{\nu_{i}^{c}}+\Delta_{(2+i)(7+i)},$
-$M_{(\tilde{\nu}_{i}^{c})^{\Re}(\tilde{\nu}_{j}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=\Delta_{(2+i)(7+i)}.$ (104)
-$\Delta_{16}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{|\mu|^{2}\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (105)
-×$\times{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
-$\Delta_{17}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{|\mu|^{2}\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (106)
-×$\times{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\tan\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
-$\Delta_{26}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (107)
-×$\times\Bigg{\{}\ln{m_{\tilde{t}_{1}}^{2}\over m_{\tilde{t}_{2}}^{2}}+{{|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
-$\Delta_{27}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (108)
-×$\times\Bigg{\{}\ln{m_{\tilde{t}_{1}}^{2}\over m_{\tilde{t}_{2}}^{2}}+{{|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
-$\Delta_{1(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{|\mu|^{2}\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (109)
-×$\times{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
-$\Delta_{2(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (110)
-×$\times\Bigg{\{}\ln{m_{\tilde{t}_{1}}^{2}\over m_{\tilde{t}_{2}}^{2}}+{{|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
-$\Delta_{(2+i)6}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (111)
-×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
-$\Delta_{(2+i)7}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (112)
-×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}},$
-$\Delta_{(2+i)(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (113)
-×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}},$
-$\Delta_{16}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}$ (114)
-×$\times\Bigg{\{}\ln{m_{\tilde{b}_{1}}^{2}\over m_{\tilde{b}_{2}}^{2}}+{{|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
-$\Delta_{17}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (115)
-×$\times\Bigg{\{}\ln{m_{\tilde{b}_{1}}^{2}\over m_{\tilde{b}_{2}}^{2}}+{{|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}},$
-$\Delta_{26}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}$ (116)
-×$\times{{|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$
-$\Delta_{27}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (117)
-×$\times{{|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
-$\Delta_{1(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (118)
-×$\times\Bigg{\{}\ln{m_{\tilde{b}_{1}}^{2}\over m_{\tilde{b}_{2}}^{2}}+{{|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$
-$\Delta_{2(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (119)
-×$\times{{|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
-$\Delta_{(2+i)6}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (120)
-×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
-$\Delta_{(2+i)7}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (121)
-×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
-$\Delta_{(2+i)(7+j)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (122)
-×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}.$
+$M_{h_{d}(\tilde{\nu}_{i}^{c})^{\text{Im}}}^{2}\!\!\!\!\!$ $=-\text{Im}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}+\Delta_{1(7+i)},$ (99)
+$M_{h_{u}(\tilde{\nu}_{i}^{c})^{\text{Im}}}^{2}\!\!\!\!\!$ $=-\text{Im}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}+\Delta_{2(7+i)},$ (100)
+$M_{(\tilde{\nu}_{i}^{c})^{\text{Re}}\sigma_{d}}^{2}\!\!\!\!\!$ $=-\text{Im}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}+\Delta_{(2+i)6},$ (101)
+$M_{(\tilde{\nu}_{i}^{c})^{\text{Re}}\sigma_{u}}^{2}\!\!\!\!\!$ $=\text{Im}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}+\Delta_{(2+i)7},$ (102)
+$M_{(\tilde{\nu}_{i}^{c})^{\text{Re}}(\tilde{\nu}_{i}^{c})^{\text{Im}}}^{2}\!\!\!\!\!$ $=2\text{Im}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}v_{u}$ (103)
+$-(A_{\kappa}\kappa)_{i}\text{Re}(e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})\upsilon_{\nu_{i}^{c}}+\Delta_{(2+i)(7+i)},$
+$M_{(\tilde{\nu}_{i}^{c})^{\text{Re}}(\tilde{\nu}_{j}^{c})^{\text{Im}}}^{2}\!\!\!\!\!$ $=\Delta_{(2+i)(7+i)}.$ (104)
+$\Delta_{16}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{|\mu|^{2}\cot\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (105)
+×$\times{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
+$\Delta_{17}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{|\mu|^{2}\cot\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (106)
+×$\times{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\tan\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
+$\Delta_{26}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (107)
+×$\times\Bigg{\{}\ln{m_{\tilde{t}_{1}}^{2}\over m_{\tilde{t}_{2}}^{2}}+{{|A_{t}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
+$\Delta_{27}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (108)
+×$\times\Bigg{\{}\ln{m_{\tilde{t}_{1}}^{2}\over m_{\tilde{t}_{2}}^{2}}+{{|A_{t}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
+$\Delta_{1(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{|\mu|^{2}\cot\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (109)
+×$\times{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
+$\Delta_{2(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (110)
+×$\times\Bigg{\{}\ln{m_{\tilde{t}_{1}}^{2}\over m_{\tilde{t}_{2}}^{2}}+{{|A_{t}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
+$\Delta_{(2+i)6}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (111)
+×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
+$\Delta_{(2+i)7}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (112)
+×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}},$
+$\Delta_{(2+i)(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (113)
+×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}},$
+$\Delta_{16}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}$ (114)
+×$\times\Bigg{\{}\ln{m_{\tilde{b}_{1}}^{2}\over m_{\tilde{b}_{2}}^{2}}+{{|A_{b}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
+$\Delta_{17}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (115)
+×$\times\Bigg{\{}\ln{m_{\tilde{b}_{1}}^{2}\over m_{\tilde{b}_{2}}^{2}}+{{|A_{b}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}},$
+$\Delta_{26}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}$ (116)
+×$\times{{|\mu|^{2}\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$
+$\Delta_{27}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (117)
+×$\times{{|\mu|^{2}\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
+$\Delta_{1(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (118)
+×$\times\Bigg{\{}\ln{m_{\tilde{b}_{1}}^{2}\over m_{\tilde{b}_{2}}^{2}}+{{|A_{b}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$
+$\Delta_{2(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (119)
+×$\times{{|\mu|^{2}\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
+$\Delta_{(2+i)6}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (120)
+×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
+$\Delta_{(2+i)7}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (121)
+×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
+$\Delta_{(2+i)(7+j)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (122)
+×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}.$